### v1

# Selective labeling with false discovery rate control

Abstract
Obtaining high-quality labels for large datasets is expensive, requiring massive annotations from human experts. While AI models offer a cost-effective alternative by predicting labels, their label quality is compromised by the unavoidable labeling errors. Existing methods mitigate this issue through selective labeling, where AI labels a subset and human labels the remainder. However, these methods lack theoretical guarantees on the quality of AI-assigned labels, often resulting in unacceptably high labeling error within the AI-labeled subset. To address this, we introduce Conformal Labeling, a novel method to identify instances where AI predictions can be provably trusted. This is achieved by controlling the false discovery rate (FDR), the proportion of incorrect labels within the selected subset. In particular, we construct a conformal $p$-value for each test instance by comparing AI models’ predicted confidence to those of calibration instances mislabeled by AI models. Then, we select test instances whose $p$-values are below a data-dependent threshold, certifying AI models’ predictions as trustworthy. We provide theoretical guarantees that Conformal Labeling controls the FDR below the nominal level, ensuring that a predefined fraction of AI-assigned labels is correct on average. Extensive experiments demonstrate that our method achieves tight FDR control with high power across various tasks, including image and text labeling, and LLM QA.
1 Introduction
Large-scale, high-quality labeled data is crucial for the machine learning pipelines (Deng et al., 2009). While experts could provide high-quality labels for moderately sized datasets, the growing size of modern datasets has made this approach prohibitively expensive. AI models offer a cost-effective alternative by predicting labels, bypassing the need for human experts. However, AI models are prone to labeling error (Northcutt et al., 2021; Tan et al., 2024). For example, empirical evidence demonstrates that even state-of-the-art LLMs exhibit high labeling error when used for text annotation Baumann et al. (2025). The labeling error inherent to AI models significantly compromises their label quality, hindering the deployment of AI labeling for production. To balance the trade-off between labeling cost and error, selective labeling has been a promising solution (Geifman & El-Yaniv, 2017; Wang et al., 2023) by combining AI predictions with expert annotations.
Prior work on selective labeling primarily designed heuristic methods (Wang et al., 2021; Bernhardt et al., 2022; Wang et al., 2024a) that rely on AI models to label high-confidence instances while deferring the rest to human experts. However, these methods do not provide any theoretical guarantees on label quality. To address this, a recent work (Candès et al., 2025) proposes probably approximately correct (PAC) labeling, which guarantees that the overall labeling error is controlled with high probability. They achieve the guarantee by counterbalancing the error of AI labels with the zero error of expert labels. While the guarantee is theoretically appealing, the labeling error of AI models can be unacceptably high, even reaching 100%, while the overall labeling error is controlled. These limitations motivate us to investigate how to provably guarantee the quality of AI-assigned labels.
In this work, we propose Conformal Labeling, which formulates the labeling problem as multiple hypotheses testing. This method allows us to select a subset of the unlabeled dataset for AI labeling and ensure the quality of AI-assigned labels by providing a rigorous guarantee on the false discovery rate (FDR). In particular, given a labeled calibration dataset, we compute conformal $p$-values for test instances through a rank-based comparison of their predicted confidence against those of calibration instances where AI predictions are incorrect. This ensures that $p$-values for mislabeled instances stochastically dominate the uniform distribution on [0, 1], while $p$-values for correctly labeled instances are concentrated near zero, allowing selection procedures to distinguish between them easily. Then, we select test instances whose $p$-values are less than or equal to a data-driven threshold, which is delicately set with the calibration dataset to control FDR at the desired level. We further theoretically prove that Conformal Labeling achieves valid FDR control under mild assumptions, ensuring that the expected proportion of incorrect labels in the selected subset is below the desired level. While the label quality of existing methods depends heavily on model performance, Conformal Labeling guarantees label quality regardless of the underlying model’s performance by strictly controlling the FDR.
To validate our method, we conduct extensive experiments on various labeling tasks, including image labeling (ImageNet (Deng et al., 2009), ImageNet-V2 (Recht et al., 2019)), text labeling (stance on global warming (Luo et al., 2021), and misinformation (Gabriel et al., 2022)), and LLM QA (MedMCQA (Pal et al., 2022), MMLU (Hendrycks et al., 2021), MMLU-Pro (Wang et al., 2024a)) tasks. The results demonstrate that Conformal Labeling achieves high power with controlled FDR, indicating that AI models can label a large proportion of data with high quality. For example, Conformal Labeling can label 58.67% of the ImageNet dataset with ResNet-34 (He et al., 2016), while keeping the FDR below 10%. In comparison, a naive approach of using AI-assigned labels for the entire dataset results in a labeling error of over 25%. Moreover, through comprehensive ablation studies, we validate that Conformal Labeling is robust to the size of calibration datasets, and more powerful models enable better selection results.
We summarize our contributions as follows:
  • We propose Conformal Labeling, a novel method for identifying instances where AI predictions could be provably trusted. Regardless of AI models’ performance, Conformal Labeling guarantees the quality of AI-assigned labels by strictly controlling the FDR.
  • We theoretically prove that Conformal Labeling provides a strict quality guarantee for AI-assigned labels: it achieves valid FDR control, ensuring the expected proportion of incorrect labels is below a user-specified level.
  • We empirically show that Conformal Labeling significantly reduces the labeling cost while tightly controlling the FDR, through extensive experiments conducted on image labeling, text labeling, and LLM QA tasks with various models.
2 Preliminaries
Problem setup.
In this work, we study the problem of identifying instances where AI predictions can be provably trusted. Here, we give a formulation of multi-class classification as an example. Let $\mathcal{X}$ denote the feature space and $\mathcal{Y}=\{1,\ldots,K\}$ denote the label space. The test dataset $\mathcal{D}_{\mathrm{test}}=\{X_{j}\}_{j=1}^{m}$ consists of $m$ data instances, sampled i.i.d. from a data distribution $\mathcal{P}_{\mathcal{X}}$. We denote the unseen ground-truth labels of instance $X_{j}$ as $Y_{j}$. Besides, we consider a pre-trained AI model $f:\mathcal{X}\to\mathbb{R}^{|\mathcal{Y}|}$ used to generate labels for the test dataset. For a given test instance $X$, the AI model predicts the label with the largest estimated probability $\hat{Y}=\arg\max_{y\in\mathcal{Y}}f_{y}(X)$, where $f_{y}(X)$ denotes the estimated class probability for class $y\in\mathcal{Y}$.
Since AI models are typically prone to labeling errors, we aim to select a large subset from the test dataset $\mathcal{D}_{\mathrm{label}}$ to control the portion of incorrect labels. Formally, our goal is to identify the largest subset of indices $\mathcal{R}\subseteq\{1,\cdots,m\}$ that controls the false discovery rate (FDR), defined as below:
$\text{FDR}=\mathbb{E}\left[\frac{|\mathcal{R}\cap\mathcal{H}_{0}|}{\max(|\mathcal{R}|,1)}\right],$
where $H_{0}=\{j\in\{1,\cdots,m\}:Y_{j}\neq\hat{Y}_{j}\}$ is the set of indices with incorrect predictions. For notation shorthand, we denote $m]=\{1,\cdots,m\}$ in the following. The FDR metric measures the expected proportion of mislabeled samples within the AI-labeled subset, illustrating the quality of AI-assigned labels by explicitly bounding the fraction of incorrect labels.
In addition to FDR control, we also expect AI models to label as many test instances as possible, which corresponds to maximizing power:
$\text{Power}=\mathbb{E}\left[\frac{|\mathcal{R}\cap\mathcal{H}_{1}|}{\max(|\mathcal{H}_{1}|,1)}\right],$ (1)
[…]
In this work, we assume access to a small labeled calibration dataset $\mathcal{D}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$. For convenience, we denote the test dataset as $\mathcal{D}_{\mathrm{test}}=\{(X_{j},Y_{j})\}_{j=n+1}^{n+m}$, where $Y_{j}$ is not observed. Since the labeling cost of a small dataset by human annotators is typically affordable, this assumption is practical in the real world and is also adopted in prior work (Candès et al., 2025). Besides, we assume that examples of the test and calibration datasets are both drawn i.i.d. from the joint distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$, a common setting in selective labeling (Jung et al., 2024; Candès et al., 2025).
Selective labeling methods and their limitations.
Ensuring high-quality labels while reducing annotation costs has motivated extensive research on selective labeling. Prior work on selective labeling primarily focused on heuristic methods. For example, some studies design collaborative annotation frameworks that combine expert labels with LLM labels to streamline the annotation process (Li et al., 2023; Kim et al., 2024). Others propose domain-specific methods, such as meta-learning strategies for medical image labeling (Vrabac et al., 2022), annotation frameworks for text data (Duan & Lalor, 2023), and human–AI collaborative systems for object detection (Zhang et al., 2025). Although these heuristic approaches effectively reduce annotation costs, they lack formal guarantees on the label quality, which can result in unreliable labels when AI models perform poorly.
Probably approximately correct (PAC) labeling (Candès et al., 2025) addresses this limitation by providing a theoretical guarantee: the overall labeling error across the dataset is controlled with high probability. At its core, PAC labeling strategically collects zero-error expert labels for instances where the AI model exhibits the highest uncertainty, while relying on potentially imperfect AI predictions for more certain instances. This strategic allocation ensures that the dataset’s overall labeling error is small, as the zero-error expert annotations effectively counterbalance the error introduced by the AI-assigned labels. However, because the guarantee only applies at the aggregate dataset level, the subset labeled by AI may exhibit high labeling error. This limitation underscores a critical gap: existing selective labeling methods cannot ensure the quality of AI-assigned labels, hindering their reliable deployment in real-world applications. To address this, we explore methods to provably guarantee the quality of AI-assigned labels in selective labeling.
3 Method
3.1 Conformal Labeling
Our previous section shows that existing methods cannot guarantee the quality of AI-assigned labels. To address this limitation, we propose Conformal Labeling, which identifies instances where AI predictions can be provably trusted by controlling the FDR. Our approach is composed of three primary steps: quantifying uncertainty, constructing conformal $p$-values, and thresholding.
Uncertainty quantification.
Our approach builds on a key insight: we should select instances where the model exhibits high confidence in its predictions. To quantify the model confidence, we define an uncertainty score $\mathcal{S}:\mathcal{X}\to\mathbb{R}$, where a higher value indicates greater model uncertainty. We note that in the conformal inference framework, this score function is also known as the non-conformity score function. For example, we employ $\mathcal{S}(X)=1-\max_{y\in\mathcal{Y}}f_{y}(X)$ as our uncertainty score function (Hendrycks & Gimpel, 2016). This score is a valid measure of uncertainty, since prior works establish that misclassified samples generally receive lower probability scores (i.e., $\max_{y\in\mathcal{Y}}f_{y}(X)$) than correctly classified ones (Hendrycks & Gimpel, 2016; Tu et al., 2024).
Statistical guarantee via conformal $p$-value.
To provide a statistical guarantee, we reformulate our problem as the following multiple hypothesis testing problem:
$H_{j}^{0}:Y_{n+j}\neq\hat{Y}_{n+j}\quad\text{v.s.}\quad H_{j}^{1}:Y_{n+j}=\hat{Y}_{n+j},\quad\forall j=1,\cdots,m$
Rejecting the null hypothesis $H_{j}^{0}$ indicates that $(X_{n+j},\hat{Y}_{n+j})$ should be included in the subset, as it is deemed to be classified correctly. To construct the selection subset, we employ conformal p-value, which builds upon conformal inference framework (Vovk et al., 1999; 2005). The underlying intuition is that: a test instance $X$ is likely a misclassification if its uncertainty score $\mathcal{S}(X)$ is generally larger than the scores of instances that are known to be misclassified. Leveraging this idea, conformal $p$-value is computed through a rank-based comparison of $\mathcal{S}(X)$ against uncertainty scores of misclassified instances.
Algorithm 1 Conformal Labeling
1:Mislabeled Calibration set $\mathcal{D}^{0}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n_{0}}$, test instances $\{X_{n+j}\}_{j=1}^{m}$, pre-trained classifier $f$, calibration set size $n=|\mathcal{D}_{\mathrm{cal}}|$, FDR target $\alpha\in(0,1)$.
2:# 1. Compute uncertainty scores
3:for $i=1,\dots,n+m$ do
4: Compute $\mathcal{S}_{i}:=1-\max_{y\in\mathcal{Y}}f_{y}(X_{i})$.
5:end for
6:# 2. Construct conformal p-values
7:for $j=1,\dots,m$ do
8: Construct $\hat{p}_{j}$ according to equation 2.
9:end for
10:# 3. Thresholding
11:Compute $j^{*}=\max\left\{j:\hat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\}$, where $\hat{p}_{(j)}$ is the $j$-th smallest p-value.
12:Output: $\mathcal{R}=\{j:\hat{p}_{j}\leq\hat{p}_{(j^{*})}\}$.
Formally, for the calibration dataset $\mathcal{D}_{cal}$, we identify the subset $\mathcal{D}_{cal}^{0}\subseteq\mathcal{D}_{cal}$ where instances are misclassified by the AI model. For simplicity, we denote $\mathcal{D}_{cal}^{0}=\{(X_{i},Y_{i})\}_{i=1}^{n_{0}}$, and thus $Y_{i}\neq\hat{Y}_{i}$ for $i=1,\cdots,n_{0}$. We compute the uncertainty scores for the entire dataset $\{(X_{i},Y_{i})\}_{i=1}^{n+m}$: $\mathcal{S}_{i}=1-\max_{y\in\mathcal{Y}}f_{y}(X_{i})$. Then, the conformal $p$-value for the instance $X_{n+j}$ is computed by
$\hat{p}_{j}=\frac{\sum_{i=1}^{n_{0}}\mathbf{1}\{\mathcal{S}_{i}<\mathcal{S}_{n+j}\}+\left(1+\sum_{i=1}^{n_{0}}\mathbf{1}\{\mathcal{S}_{i}=\mathcal{S}_{n+j}\}\right)\cdot U_{j}}{n_{0}+1},$ (2)
where $U_{j}\sim\mathrm{Uniform}[0,1]$ are i.i.d. uniform random variable to randomize over ties when $S_{n+j}$ equals some $S_{i}$, ensuring a continuous conformal $p$-value. Here, $\hat{p}_{j}$ quantifies how extreme the uncertainty of $X_{n+j}$ is compared to the scores of misclassified instances, with a small $\hat{p}_{j}$ providing strong statistical evidence for correct prediction.
Standard results from conformal inference establish the validity of conformal $p$-value in practice: if the instance $X_{n+j}$ is misclassified, then the conformal $p$-value stochastically dominates the uniform distribution on $[0,1]$ (Bates et al., 2023; Jin & Candès, 2023): $\mathbb{P}\{\hat{p}_{j}\leq\alpha\mid H_{j}^{0}\text{ is true}\}\leq\alpha.$ This property indicates that the conformal $p$-values for misclassified instances are biased to be high, which allows us to set a threshold to flag potential correct instances, while controlling the FDR.
Thresholding.
After obtaining conformal $p$-values for test instances, we apply a thresholding rule inspired by the Benjamini–Hochberg (BH) procedure (Benjamini & Hochberg, 1995) to select a maximal subset $\mathcal{R}$ for AI labeling while controlling the FDR at level $\alpha$. The key idea is that we gradually increase the acceptance threshold until including additional samples would risk exceeding the desired labeling error. In particular, let $p_{(1)}\leq\ldots\leq p_{(m)}$ denote the ordered statistics of the $p$-values; the rejection set of our selection procedure applied to the conformal $p$-values is $\mathcal{R}=\{j:\hat{p}_{j}\leq\hat{p}_{(j^{*})}\}$, where
$j^{*}=\max\left\{j:\hat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\},$
[…]
Theoretical results.
We now provide a theoretical guarantee for Conformal Labeling. In the following theorem, we establish that Algorithm 1 controls the FDR below the desired level $\alpha$. This theorem ensures that AI predictions can be provably trusted in the selected subset, as the expected proportion of incorrect labels is controlled below $\alpha$. The proof is provided in Appendix B.2.
Theorem 3.1.
Suppose calibration samples $(X_{i},Y_{i})_{i=1}^{n}$ and test samples $(X_{n+j},Y_{n+j})_{j=1}^{m}$ are i.i.d. Let $\alpha\in(0,1)$ be the target FDR level, and suppose the selection set $\mathcal{R}$ is determined by Algorithm 1 applied at the target FDR level $\alpha$. Define $p=\mathbb{E}[H_{j}^{0}]$, the probability that a test sample $(X_{n+j},Y_{n+j})$ is incorrectly predicted. Then, the FDR of the selection set $\mathcal{R}$ satisfies:
$\text{FDR}\leq\left[1-(1-p)^{n+1}\right]\alpha\leq\alpha.$
3.2 Choice of uncertainty score
In the above analysis, we establish that Conformal Labeling controls the FDR below the desired level. However, this guarantee alone is insufficient: a trivial procedure that simply labels nothing would achieve a perfect FDR of 0, yet offer no practical value. This highlights the need to also evaluate the method’s statistical power, which measures the method’s ability to identify as many correctly labeled instances as possible (see Eq. (1)).
As shown in prior work (Jin & Candès, 2023; Gui et al., 2024; Bai et al., 2025b), the statistical power of this method depends on the quality of the uncertainty score. In particular, a score that better separates correct from incorrect predictions directly increases statistical power (see Proposition 7 of Jin & Candès (2023)). To deliver practical recommendations, we empirically compare several uncertainty scores by visualizing their resulting $p$-value distributions and measuring the final statistical power employed with our method. We utilize a pre-trained ResNet-34 model on the ImageNet dataset, with three uncertainty scores: maximum softmax probability (MSP) (Hendrycks & Gimpel, 2016), energy score (Liu et al., 2020), and DOCTOR-$\alpha$ score ($D_{\alpha}$) (Granese et al., 2021). We give an overview of these score functions in Appendix C. The results in Figure 1 show that both MSP and $D_{\alpha}$ score provide a clear distinction between correct and incorrect predictions, thus achieving high statistical power. However, the energy score fails to provide this separation, leading to low power. Given the comparable power performance of MSP and $D_{\alpha}$ score, we will use the more computationally efficient MSP in our main experiments.
(a) MSP
(b) Energy
(c) $D_{\alpha}$
(d) Power comparison
Figure 1: Empirical distributions and power (employed with our method) of conformal $p$-values under different uncertainty scores. The experiments include maximum softmax probability (MSP), energy, and DOCTOR-$\alpha$ score ($D_{\alpha}$). The experiments are conducted on ImageNet with ResNet-34. The results show that both MSP and $D_{\alpha}$ score create a clear distinction between correct and incorrect predictions, thus achieving high statistical power. However, the energy score fails to provide this separation, leading to low power.
4 Experiments
In this section, we evaluate the effectiveness of Conformal Labeling on image labeling, text labeling, and LLM QA tasks with various models. We find that it achieves tight FDR control and high power, indicating that AI models can label a large proportion of data with high quality. We also conduct comprehensive ablation studies to provide practical guidance for applying our method.
4.1 Experimental setup
Tasks and datasets.
We evaluate the effectiveness of Conformal Labeling across three labeling tasks, including image labeling, text labeling, and LLM question answering (QA). In appendix G, we also demonstrate how to extend Conformal Labeling to regression tasks. For the LLM QA task, the goal is to identify subsets of questions that LLMs can answer correctly. We employ common benchmark datasets for evaluations in each labeling task. For image classification, we use ImageNet (Deng et al., 2009) and its variant ImageNet-V2 datasets (Recht et al., 2019). For text labeling, we adopt two benchmark datasets. The first is Stance on Global Warming (Luo et al., 2021), which provides annotations ($Y_{i}\in\{\text{agree},\text{neutral},\text{disagree}\}$) to judge whether a headline agrees that global warming is a serious concern. The second is Misinformation (Gabriel et al., 2022), which contains binary annotations ($Y_{i}\in\{\text{misinfo},\text{real}\}$) for identifying whether a given text contains misinformation. For the LLM QA task, we evaluate our method on MedMCQA (Pal et al., 2022), MMLU (Hendrycks et al., 2021), and MMLU-Pro (Wang et al., 2024b) datasets.
Table 1: Performance of Conformal Labeling on three labeling tasks. $\uparrow$ indicates that a larger value is better. We evaluate on Image Labeling (ImageNet, ImageNet-V2), text Labeling (Stance on global warming, Misinformation), and LLM QA (MedMCQA, MMLU) tasks. We report results for Conformal Labeling at $\alpha=0.1$ and compare Conformal Labeling against two baselines: (i) a naive approach of labeling instances whose uncertainty score $\mathcal{S}_{n+j}\leq 0.1$ with AI models, and (ii) labeling the entire dataset with AI models. The results show that Conformal Labeling consistently achieves tighter FDR control across all datasets and models compared with the baseline.
Task Dataset Model Conformal Labeling ($\alpha$= 0.1) Naive ($\mathcal{S}\leq 0.1$) AI only
FDR % Power %($\uparrow$) Ratio %($\uparrow$) FDR % Power %($\uparrow$) Ratio %($\uparrow$) Error %($\downarrow$)
Image ImageNet ResNet-34 9.97 80.01 58.67 4.79 63.09 43.71 26.71
DenseNet-161 9.99 85.03 65.56 5.58 72.08 52.98 22.89
ResNeXt50 10.00 86.06 66.83 6.08 75.57 56.21 22.39
CLIP-VIT-B/32 9.98 46.04 27.47 5.53 28.65 16.28 40.35
ImageNet-V2 ResNet-34 10.00 61.99 37.87 7.25 56.03 33.15 39.03
DenseNet-161 9.83 66.67 43.39 9.39 65.63 42.45 34.88
ResNeXt50 9.95 67.86 44.75 10.17 68.71 45.38 34.08
CLIP-VIT-B/32 9.96 34.56 18.10 7.86 25.83 13.18 47.78
Text Stance Llama-3.1-8B-Instruct 9.77 19.02 9.42 24.43 54.94 31.39 52.04
Qwen3-32B 9.56 10.70 7.25 14.42 16.64 11.11 36.52
Qwen2.5-72B-Instruct 9.71 26.97 17.84 26.27 74.49 59.01 35.09
Misinformation Llama-3.1-8B-Instruct 9.88 7.31 5.81 18.38 66.17 55.25 24.28
Qwen3-32B 9.91 49.21 37.48 10.77 52.27 40.03 24.08
Qwen2.5-72B-Instruct 9.58 44.36 34.81 17.81 87.13 74.72 21.69
QA MedMCQA Llama-3.1-8B-Instruct 9.70 31.44 18.90 15.01 46.74 29.53 40.35
Qwen3-32B 9.75 49.80 33.27 13.79 65.27 45.36 33.44
Llama-3.1-70B-Instruct 9.95 69.67 49.59 4.52 48.92 32.79 28.90
MMLU Llama-3.1-8B-Instruct 9.99 58.25 37.47 8.47 53.72 33.96 35.72
Qwen3-32B 10.00 82.96 65.22 8.13 78.47 60.40 21.43
Llama-3.1-70B-Instruct 9.96 88.20 72.10 4.18 67.94 52.17 18.24
Models.
We conduct extensive experiments on various open-sourced AI models. For image classification, we utilize three well-established deep image classifiers: ResNet-34 (He et al., 2016), DenseNet-161 (Huang et al., 2018), and ResNext50 (Xie et al., 2017). Additionally, we employ the Vision-Language Model CLIP (Radford et al., 2021), which is based on a Vision Transformer architecture (ViT-B/32) (Dosovitskiy et al., 2021). The above classifiers are provided by TorchVision (Paszke et al., 2019). For text labeling, we employ three LLMs: Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Qwen3-32B, Qwen2.5-72B-Instruct (Qwen et al., 2025). For LLM QA tasks, we employ five LLMs: Qwen3-8B (Yang et al., 2025), Qwen3-14B, Qwen3-32B, Llama-3.1-8B-Instruct, and Llama-3.1-70B-Instruct. The above LLMs are provided by Hugging Face.
Baselines and evaluation metrics.
We evaluate Conformal Labeling with two baseline methods: using AI models to label test instances with uncertainty scores $\mathcal{S}_{n+j}\leq 0.1$, and applying AI predictions to the entire test dataset. We compare Conformal Labeling’s selection procedure against BH (Benjamini & Hochberg, 1995), Storey-BH (Storey, 2002), and Quantile-BH procedures (Benjamini et al., 2006). We evaluate the performance of our method and baseline using the following metrics: (1) FDR, the expected proportion of incorrect labels in the selected set $\mathcal{R}$; (2) Power, the proportion of correctly labeled instances selected; (3) AI-labeled ratio, the number of data labeled by the AI models divided by the combined size of the calibration and test datasets.
Implementation details.
To ensure the reliability of our results, we repeat each experiment 1000 times and report the average result. We randomly select 10% of the data as the calibration dataset. For all the experiments, we use the maximum softmax probability (MSP) as the uncertainty score function. In the LLM QA task, we adopt the standard multiple-choice evaluation pipeline: given a question and candidate answers (e.g., A, B, C, or D), the model estimates the probability of each option by extracting the logits corresponding to the option tokens and applying a softmax transformation, with the predicted label taken as the option with the highest probability. For text labeling, we reformulate each sample into a multiple-choice format (e.g., “positive,” “negative,” or “neutral”), enabling the same probability-extraction procedure as in the LLM QA task. More details of implementation are provided in Appendix F. We provide the code for reproducing our main experiments in this anonymous repository.
4.2 Experiments results
Conformal Labeling achieves tight FDR control with high power.
In table 1, we present the performance of Conformal Labeling against two baselines on three different labeling tasks: image labeling, text labeling, and LLM QA. A salient observation is that across all the labeling tasks and all the model architectures, Conformal Labeling successfully controls the FDR at or below the target FDR level. In comparison, both baseline methods lack FDR control, resulting in substantial labeling errors that compromise label quality when AI models are inaccurate. It is worth noting that the FDR is tightly controlled: for $\alpha=10\%$, most experiments yield FDRs below $9.9\%$, with the largest deviation at $9.56\%$. This tight FDR control directly leads to high selection power. For example, on MMLU with Qwen3-32B at $\alpha=10\%$, Conformal Labeling achieves a power of $82.96\%$, labeling $65.22\%$ of the dataset with the AI model. Overall, empirical results show that Conformal Labeling consistently achieves tight FDR control with high power across different datasets and models.
Figure 2: Performance comparison of Conformal Labeling across models of varying accuracy. We employ Qwen3-8B, Qwen3-14B, and Qwen3-32B (model accuracy increases with parameter count) on MMLU and MMLU-Pro with $\alpha=0.1$. The results show that model with higher accuracy achieves greater power and AI-labeled ratio.
Higher prediction accuracy enables better selection results.
The performance of Conformal Labeling depends heavily on the underlying prediction accuracy, which is influenced by model capacity and dataset difficulty. We evaluate Conformal Labeling with Qwen3-8B, Qwen3-14B, and Qwen3-32B, whose increasing scales provide greater capacity. MMLU-Pro, more challenging than MMLU, results in lower prediction accuracy. Figure 2 presents the evaluation results. In all cases, the realized FDR remains below and close to $\alpha=0.1$. Higher accuracy—achieved with stronger models or easier datasets—boosts power and the AI-labeled ratio. Specifically, for any given model, performance is superior on MMLU compared to MMLU-Pro. Similarly, for each dataset, larger models yield greater power and AI-labeled ratios. Overall, higher prediction accuracy leads to better selection results by achieving higher power and AI-labeled ratio with controlled FDR.
How many calibration samples are needed?
The size of the calibration set plays a crucial role in constructing reliable conformal $p$-values. To study the effect of $|D_{\mathrm{cal}}|$ on FDR and power, in Figure 3, we label {5%, 10%, 20%} of the unlabeled dataset as the calibration dataset. Our results demonstrate that Conformal Labeling is robust to calibration set size: even with a 5% calibration ratio, the FDR remains controlled with low standard deviation. Increasing the calibration ratio reduces the variance of both FDR and power, although the improvement from 10% to 20% is negligible. Based on this trade-off between variance reduction and labeling cost, we use a 10% calibration ratio for all the experiments. In summary, while Conformal Labeling is robust to calibration set sizes, larger calibration sets reduce the variance of FDR and power, thereby enhancing selection stability.
Figure 3: Performance of Conformal Labeling with varying calibration set sizes on the MedMCQA dataset. The top row corresponds to the results from Qwen3-8B and the bottom row to those from Qwen3-32B; each column corresponds to a value of calibration ratio. Shaded regions indicate one standard deviation around the average. The results show that a large calibration set slightly reduces the variance of FDR and power, leading to a more robust selection outcome. Overall, our method is robust to changes in the calibration set size.
Figure 4: Albation study on the selection procedure. The ablation study is conducted on the MMLU dataset with Qwen3-32B. We substitute Conformal Labeling’s selection procedure with alternative selection procedures, including BH, Storey-BH, and Quantile-BH procedures. The figure shows the Conformal Labeling’s selection procedure consistently achieves the tightest FDR control and the highest power among all the selection procedures.
Ablation study on the selection procedure.
To validate the effectiveness of Conformal Labeling’s selection procedure, we conduct an ablation study substituting it with alternative selection procedures, including BH, Storey-BH (Storey, 2002), and Quantile-BH (Benjamini et al., 2006) procedures (see Appendix D for details). The hyperparameters for the Storey-BH and Quantile-BH procedures are chosen using a bootstrap method detailed in the Appendix E. An ideal selection procedure aims to maximize power while keeping the FDR below the desired level.
The results of this ablation study are presented in Figure 4. While all the selection procedures successfully control the FDR below the desired level, Conformal Labeling’s selection procedure (red line) provides the tightest control, with realized FDR consistently closest to the desired level (dashed blue line). In comparison, alternative selection procedures are more conservative, especially at low target FDR levels, which reduces power. Overall, this ablation study highlights the superiority of Conformal Labeling’s selection procedure.
5 Related work
Selection with conformal $p$-values.
Conformal $p$-values are widely studied for their distribution-free, model-agnostic properties in selection tasks. Existing methods fall into two main categories: conformal novelty detection (Bates et al., 2023; Bashari et al., 2023; Wu et al., 2025; Bashari et al., 2025; Lee et al., 2025; Huo et al., 2025) and conformal selection (Jin & Candès, 2023; Gui et al., 2024). Conformal novelty detection aims to identify out-of-distribution instances, while conformal selection aims to select data points that meet a specific quality criterion. For conformal novelty detection, recent advances primarily focus on improving the power of the selection procedure by using various forms of side information (Liang et al., 2024; Marandon et al., 2024; Zhao & Sun, 2025). Jin & Candès (2023) introduced the conformal selection framework for data selection in regression settings, which has since been extended to various settings: multivariate data selection (Bai et al., 2025b), online data selection (Xu & Ramdas, 2024; Liu et al., 2025), and human-in-the-loop adaptive data selection (Gui et al., 2025). Recent works have also applied CS to different tasks such as candidate screening (Lu et al., 2025), drug discovery (Jin & Candès, 2023; Bai et al., 2025a), and foundation model alignment (Gui et al., 2024). However, none of them has studied how to construct conformal $p$-values and design selection procedures in the selective labeling task.
Selective labeling.
Our method builds on the idea of selective labeling, which prioritizes the collection of expert labels for instances where the AI model exhibits uncertainty, and relies on the model’s prediction where it is confident. (Gu et al., 2012; Vrabac et al., 2022). Prior works on selective labeling have explored various heuristic methods, including collaborative annotation frameworks that integrate expert and LLM efforts (Li et al., 2023), and domain-specific applications in text, vision, and medicine (Kim et al., 2024; Vrabac et al., 2022; Duan & Lalor, 2023; Zhang et al., 2025). More recent research aims to provide theoretical foundations, with methods designed to control the overall labeling error (Candès et al., 2025) or to enable valid statistical inference (Zrnic & Candes, 2024; Gligorić et al., 2024). Our method is also closely related to selective prediction, where models are allowed to abstain from making a prediction when uncertain about the output (Geifman & El-Yaniv, 2017; Mozannar & Sontag, 2020; Yang et al., 2023). For LLMs, recent works have also studied how to teach the model not to predict when the model is uncertain (Kamath et al., 2020; Yoshikawa & Okazaki, 2023). These selective prediction methods also lack theoretical guarantees for the AI predictions. Distinct from prior methods, our work is the first to provide a guarantee for the quality of AI-assigned labels, thereby enabling the reliable deployment of selective labeling.
6 Conclusion
In this work, we propose Conformal Labeling, a novel selective labeling method for identifying samples where AI predictions can be provably trusted. This is achieved by controlling the false discovery rate (FDR), which ensures that the expected fraction of incorrect labels in the selected subset is below a user-specified level. The key idea is to reformulate selective labeling as multiple hypothesis testing, which enables distinct theoretical guarantees and methodological advantages compared to prior approaches. In particular, we construct a conformal $p$-value for each test instance by comparing the AI model’s predicted confidence to those of mislabeled calibration instances. Then, we select all the test samples whose $p$-values are smaller than or equal to a data-dependent threshold. We theoretically prove that Conformal Labeling successfully controls the FDR under mild assumptions. Extensive experiments demonstrate that Conformal Labeling achieves tight FDR control and high power across various tasks, including image and text labeling, and LLM QA. We hope the insight from this work will inspire future research to explore reliable selective labeling.
Appendix A Use of Large language model
This paper uses large language models solely to polish specific sentences or paragraphs, without further use of LLMs for other purposes.
Appendix B Technical proofs
B.1 A lemma for proving Theorem 3.1
Lemma B.1.
Under the assumptions of Theorem 3.1, where the calibration samples $(X_{i},Y_{i})_{i=1}^{n}$ and the test samples $(X_{n+j},Y_{n+j})_{j=1}^{m}$ are independently and identically distributed (i.i.d.), and conditional on $n_{0}$ (the number of true null hypotheses in the calibration set) and $m_{0}$ (the number of true null hypotheses in the test set), the expected false discovery proportion (FDP) satisfies
$\mathbb{E}[\text{FDP}\mid n_{0},m_{0}]\leq\frac{1+n}{1+n_{0}}\frac{m_{0}}{m}\alpha,$
where $\alpha\in(0,1)$ is the target false discovery rate (FDR) level, and the selection set $\mathcal{R}$ is determined by Algorithm 1.
Proof.
Suppose the first $n_{0}$ are true null samples in the calibration set and the first $m_{0}$ samples are true null samples in the test set. By the standard result in conformal inference (Vovk et al., 2005), we have
$\mathbb{P}\{\hat{p}_{j}\leq\alpha\mid n_{0},m_{0}\}\leq\alpha,j=1,\ldots,m_{0}$
Let $j^{*}$ denote the number of rejections from Algorithm 1, and recall the false discovery proportion (FDP) as
$\text{FDP}=\frac{\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}}{\max(j^{*},1)}.$
Since $\text{FDP}=0$ whenever $j^{*}=0$, we have
$\mathbb{E}[\text{FDP}\mid n_{0},m_{0}]=$ $\mathbb{E}\!\left[\frac{1}{j^{*}}\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\middle|n_{0},m_{0},j^{*}>0\right]\mathbb{P}\left\{j^{*}>0\right\}$ (3)
$\quad\quad+\mathbb{E}\!\left[\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\middle|n_{0},m_{0},j^{*}=0\right]\mathbb{P}\left\{j^{*}=0\right\}$
$\leq$ $\mathbb{E}\!\left[\frac{1}{j^{*}}\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\middle|n_{0},m_{0},j^{*}>0\right]$
$=$ $\mathbb{E}\!\left[\text{FDP}\;\middle|\;n_{0},m_{0},j^{*}>0\right].$
By the super-uniform property of conformal $p$-values,
$\mathbb{E}\!\left[\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\;\middle|\;n_{0},m_{0},j^{*}\right]\leq m_{0}\cdot\frac{1+n}{1+n_{0}}\cdot\frac{\alpha j^{*}}{m}.$
Thus,
$\mathbb{E}\!\left[\text{FDP}\;\middle|\;n_{0},m_{0},j^{*}>0\right]\leq\frac{1}{j^{*}}\cdot m_{0}\cdot\frac{1+n}{1+n_{0}}\cdot\frac{\alpha j^{*}}{m}=\frac{m_{0}}{m}\cdot\frac{1+n}{1+n_{0}}\alpha.$
Combining with (3) yields
$\mathbb{E}[\text{FDP}\mid n_{0},m_{0}]\leq\mathbb{E}\!\left[\text{FDP}\;\middle|\;n_{0},m_{0},j^{*}>0\right]\leq\frac{m_{0}}{m}\cdot\frac{1+n}{1+n_{0}}\alpha,$
which completes the proof. ∎
B.2 Proof of theorem 3.1
Proof of theorem 3.1.
Under the i.i.d. assumption of Theorem 3.1, the calibration samples $(X_{i},Y_{i})_{i=1}^{n}$ and the test samples $(X_{n+j},Y_{n+j})_{j=1}^{m}$ are independently and identically distributed. This implies that each hypothesis, whether from the calibration or test set, has an equal probability $p$ of being a true null, where $p$ represents the expected probability of incorrect prediction under the null hypothesis.
Consequently, the number of true null hypotheses in the calibration set, denoted $n_{0}$, follows a binomial distribution $n_{0}\sim\text{Binomial}(n,p)$, and the number of true null hypotheses in the test set, denoted $m_{0}$, follows $m_{0}\sim\text{Binomial}(m,p)$. The independence across the calibration and test sets arises from the i.i.d. structure, ensuring that $n_{0}$ and $m_{0}$ are independent random variables.
Using the law of total expectation, we express FDR as
$\text{FDR}=$ $\mathbb{E}[\text{FDP}]$ (4)
$=$ $\mathbb{E}[\mathbb{E}[\text{FDP}\mid n_{0},m_{0}]]$
[…]
$=$ $\mathbb{E}\left[\frac{m_{0}}{m}\right]\cdot\mathbb{E}\left[\frac{1+n}{1+n_{0}}\right]\cdot\alpha\quad\text{since }n_{0}\text{ and }m_{0}\text{ are independent}$
$=$ $p\cdot\mathbb{E}\left[\frac{1+n}{1+n_{0}}\right]\cdot\alpha\quad\text{since }m_{0}\sim\text{Binomial}(m,p)$
Now it suffices to show that $p\cdot\mathbb{E}\left[\frac{1+n}{1+n_{0}}\right]=[1-(1-p)^{n+1}]$.
Since $n_{0}$ has probability mass function
$\mathbb{P}(n_{0}=k)=\binom{n}{k}p^{k}(1-p)^{n-k},\quad k=0,1,\dots,n,$
we compute
$p\cdot\mathbb{E}\left[\frac{1+n}{1+k}\right]=$ $p\cdot\sum_{k=0}^{n}\frac{1+n}{1+k}\binom{n}{k}p^{k}(1-p)^{n-k}$ (5)
$=$ $\sum_{k=0}^{n}\frac{(n+1)!}{(k+1)!(n-k)!}\,p^{k+1}(1-p)^{n-k}$
$=$ $\sum_{k=0}^{n}\binom{n+1}{k+1}\,p^{k+1}(1-p)^{n-k}$
$=$ $\sum_{l=1}^{n+1}\binom{n+1}{l}p^{j}(1-p)^{n+1-l}\quad\text{by letting }l=k+1$
$=$ $\sum_{l=0}^{n+1}\binom{n+1}{l}p^{j}(1-p)^{n+1-l}-\binom{n+1}{0}p^{0}(1-p)^{n+1-0}$
$=$ $\sum_{l=0}^{n+1}\mathbb{P}(X=l)-(1-p)^{n+1}\quad\text{where }X\sim\text{Binomial}(n+1,p)$
$=$ $1-(1-p)^{n+1}$
This completes the proof, establishing the desired bound on the FDR.
Appendix C Overview of Different Uncertainty Score Functions
In this section, we provide an overview of three representative uncertainty score functions that are widely used in misclassification detection: Maximum Softmax Probability (MSP) (Hendrycks & Gimpel, 2016), the energy-based score (Liu et al., 2020), and the DOCTOR-$\alpha$ score (Granese et al., 2021). Each of these functions captures predictive uncertainty from a different perspective.
Maximum Softmax Probability (MSP).
The Maximum Softmax Probability (MSP) baseline (Hendrycks & Gimpel, 2016) proposes to use confidence of the AI model as an uncertainty score.
$S_{\mathrm{MSP}}(x)=1-\max_{y\in\mathcal{Y}}p_{y}(x),$
where $p_{y}(x)$ is the softmax probability assigned to class $y$ for input $x$. The key idea is straightforward: if the model assigns a high probability to its most likely class, the prediction is considered confident. Although MSP is simple and effective, it has notable limitations: it only reflects the confidence in the top-1 prediction and ignores the structure of the remaining probability distribution, which may contain useful information about uncertainty.
Energy-based score.
The energy-based score (Liu et al., 2020) is defined as
$S_{\mathrm{Energy}}(x)=\log\sum_{y\in\mathcal{Y}}\exp\!\left(f_{y}(x)\right),$
where $f_{y}(x)$ denotes the logit value for class $y$ and $T>0$ is a temperature parameter. This score is derived from the concept of energy in statistical physics and leverages the log-sum-exp operator over all logits. Unlike MSP, which only considers the maximum probability, the energy score integrates information from the full logit vector, thereby providing a smoother and more informative confidence measure.
DOCTOR-$\alpha$ score.
The DOCTOR-$\alpha$ score (Granese et al., 2021) is defined as
$S_{\alpha}(x)=\sum_{y\in\mathcal{Y}}p_{y}(x)^{2},$
where $p_{y}(x)$ denotes the softmax probability for class $y$. This score is inspired by information-theoretic measures of uncertainty, as it is closely related to the quadratic Rényi entropy. The intuition is that if the predictive distribution is sharp (i.e., one class has probability close to one), then $\sum_{y}p_{y}(x)^{2}$ will be large, indicating high confidence. Conversely, if the distribution is flat (i.e., the model is uncertain and spreads probability mass across many classes), then $S_{\alpha}(x)$ will be small. Compared to MSP, the DOCTOR-$\alpha$ score leverages information from the entire probability distribution rather than only the top prediction, making it a richer measure of uncertainty for misclassification detection.
Appendix D BH Procedure and its Adaptive Variants
Consider testing $m$ null hypotheses $H_{0}^{1},\ldots,H_{0}^{m}$ based on their corresponding $p$-values $\{p_{1},p_{2},\ldots,p_{m}\}$. For a true null hypothesis $H_{0}^{j}$, the corresponding $p$-value $p_{j}$ is a random variable that is super-uniform on $[0,1]$ under the null hypothesis. Formally, for any $u\in[0,1]$,
$\mathbb{P}(p_{j}\leq u\mid H_{0}^{j}\text{ is true})\leq u.$
Define $p_{(j)}$ as the $j$-th smallest $p$-value among a set of $p$-values $\{p_{1},p_{2},\ldots,p_{m}\}$. Given a set of $p$-values $\{p_{1},p_{2},\ldots,p_{m}\}$, the BH algorithm returns $S=\{j\in\{0,\ldots,m\}:p_{j}\leq\frac{\alpha j^{*}}{m}\}$, where $\alpha$ is the target FDR level and
$j^{*}=\text{max}\{j\in\{1,\ldots,m\}:p_{(j)}\leq\frac{\alpha j}{m}\}$
When the null $p$-values $\{p_{j}:j\in\mathcal{H}_{0}\}$ are independent, the BH procedure is proved to control the FDR at level $\pi_{0}\alpha$ in finite samples (Benjamini & Hochberg, 1995), where $\pi_{0}=\frac{|\mathcal{H}_{0}|}{m}$ is the proportion of true nulls. The independence assumption can be further relaxed to the PRDS condition (Benjamini & Yekutieli, 2001).
If $\pi_{0}$ is small, the FDR control will be overly conservative. When $\pi_{0}$ is known, we can apply BH procedure at level $\frac{\alpha}{\pi_{0}}$ to close the gap. In practice, $\pi_{0}$ is typically unknown. Several adaptive BH procedures attempt to address this issue by estimating $\pi_{0}$ and adjusting the target FDR level $\alpha$ accordingly. These procedures are often called the $\pi_{0}$-adaptive versions of the BH algorithm. Two most famous estimators are Storey-BH (Storey, 2002) and Quantile-BH (Benjamini et al., 2006):
$\hat{\pi}_{0}^{Storey}(\lambda)=\frac{1+\sum_{i=1}^{m}\mathbf{1}\{p_{i}\geq\lambda\}}{m(1-\lambda)},\lambda\in(0,1)$
$\hat{\pi}_{0}^{Quant}(k_{0})=\frac{m-k_{0}+1}{m(1-p_{(k_{0})})},k_{0}\in\{1,\dots,m\}$
$\lambda$ and $k_{0}$ are hyperparameters determined by users.
Appendix E Hyperparameter selection for BH adaptive variants
[…]
  1. 1.
    Define a grid $R$ for the hyperparameter, i.e. $R=\{0.1,0.2,\dots,0.9\}$ for Storey-BH.
  2. 2.
    For each $\lambda\in R$, compute:
    $\widehat{\mathrm{pFDR}}_{\lambda}(\gamma)=\frac{\widehat{\pi}_{0}(\lambda)\gamma}{\widehat{\Pr}(p\leq\gamma)\{1-(1-\gamma)^{m}\}}$ (6)
    where $\widehat{\pi}_{0}(\lambda)=\widehat{\pi}_{0}^{Storey}(\lambda)\text{ or }\widehat{\pi}_{0}^{Quant}(\lambda)$ and $\widehat{\Pr}(p\leq\gamma)$ is the empirical estimate of $\Pr(p\leq\gamma)$
  3. 3.
    Generate $B$ bootstrap replicates $\{p_{1}^{*,b},\dots,p_{m}^{*,b}\}_{b=1}^{B}$ and compute $\widehat{\mathrm{pFDR}}_{\lambda}^{*,b}(\gamma)$ for each $b$.
  4. 4.
    Estimate the MSE for each $\lambda$:
    $\widehat{\mathrm{MSE}}(\lambda)=\frac{1}{B}\sum_{b=1}^{B}\left(\widehat{\mathrm{pFDR}}_{\lambda}^{*,b}(\gamma)-\min_{\lambda^{\prime}\in R}\widehat{\mathrm{pFDR}}_{\lambda^{\prime}}(\gamma)\right)^{2}$ (7)
  5. 5.
    Select $\hat{\lambda}=\operatorname*{arg\,min}_{\lambda\in R}\widehat{\mathrm{MSE}}(\lambda)$
Appendix F Implementation details
Experiment details. We run our experiments on NVIDIA GeForce RTX 4090 and NVIDIA L40 GPU, and implement all methods by PyTorch and vLLM.
Dataset details.
For the LLM QA datasets (MedMCQA, MMLU, and MMLU-Pro), we adopt the same prompts as in Luo et al. (2025). For the other datasets, we design our own prompts following the style of LLM QA prompts, as summarized in Table 2. In text labeling and LLM QA tasks, we merge the calibration and test sets to form the unlabeled dataset, except for MedMCQA. For MedMCQA, because test labels are unavailable, we use the calibration set as the unlabeled dataset. For image labeling, we use the validation sets of ImageNet and ImageNet-V2 as the unlabeled datasets.
Table 2: Prompts for different datasets
   Dataset    Prompts    LLM QA    The following are multiple-choice questions. Give ONLY the correct option, no other words or explanation:    [Question] A: [Option 1] B: [Option 2] C: [Option 3] D: [Option 4] Answer: [Mask]        Stance on    global warming    You are given a statement about climate change. Determine the stance that a human would take towards this statement.    Respond with ONLY the letter (A, B, or C) of the correct stance. Do not include any explanation.    [Input Headline] A: agrees B: neutral C: disagrees Answer: [Mask]    Misinformation    You are a fact-checking assistant. Classify the following news headline as either real (A) or misinfo (B).    Respond with ONLY the letter A or B. Do not include any explanation.    Headline: [Input text] A: real B: misinfo Answer: [Mask]
Appendix G Extension to Regression Tasks
While our primary focus is on classification, Conformal Labeling can be naturally extended to regression settings. Consider a loss function L(Y,Y^)L($\hat{Y}$) that quantifies prediction error—for instance, the squared error $L(Y,\hat{Y})=(Y-\hat{Y})^{2}$—alongside a user-specified tolerance level $\epsilon$. For each test sample, we define the null and alternative hypotheses as:
$H_{0}^{j}:L(Y_{n+j},\hat{Y}_{n+j})>\epsilon\quad\text{versus}\quad H_{1}^{j}:L(Y_{n+j},\hat{Y}_{n+j})\leq\epsilon.$
This framework generalizes the classification setting, which corresponds to the special case where $L(Y,\hat{Y})=\mathbf{1}\{Y\neq\hat{Y}\}$ and $\epsilon=0$. To apply Conformal Labeling in regression, we need an uncertainty score that reflects the model’s predictive uncertainty. While uncertainty score functions are straightforward in classification tasks (e.g., $1-\max_{y\in\mathcal{Y}}f_{y}(X)$), regression requires alternative approaches to get an uncertainty score function. For example, when using LLMs, we can leverage their verbalized confidence or prompt them to output prediction intervals, using the interval width as the uncertainty score. The complete procedure for Conformal Labeling in the regression task is outlined in Algorithm 2.
Algorithm 2 Conformal Labeling Regression Tasks
1:Calibration set $\mathcal{D}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$; test instances $\{X_{n+j}\}_{j=1}^{m}$; pre-trained predictor $f$; loss function $L$; loss threshold $\epsilon$; target FDR level $\alpha\in(0,1)$; nonconformity scores $\{S_{i}\}_{i=1}^{n+m}$
2:Identify calibration samples that exceed the loss threshold: $\mathcal{D}_{\mathrm{cal}}^{0}=\{(X_{i},Y_{i}):L(Y_{i},\hat{Y}_{i})>\epsilon\}_{i=1}^{n}$, and set $n_{0}=|\mathcal{D}_{\mathrm{cal}}^{0}|$.
3:for $j=1,\dots,m$ do
4:  Compute the conformal $p$-value $\hat{p}_{j}$ according to equation 2.
5:end for
6:Apply the step-up procedure: $j^{*}=\max\left\{j:\hat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\}$, where $\hat{p}_{(j)}$ is the $j$-th smallest $p$-value.
7:Output: The selected set $\mathcal{S}=\{j:\hat{p}_{j}\leq\hat{p}_{(j^{*})}\}$.
We evaluate Conformal Labeling on two regression problems: sentiment analysis with GPT-4o and protein structure prediction with AlphaFold. We use the data provided by Candès et al. (2025). For sentiment analysis, we prompt GPT-4o to output a prediction interval $a_{i},b_{i}]$ for each target $Y_{i}$, set the predicted value as $\hat{Y}_{i}=(a_{i}+b_{i})/2$, and use the interval length $U_{i}=b_{i}-a_{i}$ as the uncertainty score. For protein structure prediction, we take experimentally derived structures as ground truth and AlphaFold predictions as $\hat{Y}_{i}$. Uncertainty scores are obtained from AlphaFold’s internal confidence measure, the average predicted local distance difference test (pLDDT). For both experiments, we employ the L2 loss function.
In Table 3, we report the performance of Conformal Labeling on regression tasks with $\alpha=0.1$. In all cases, Conformal Labeling consistently controls the realized FDR below the target level. A key observation is that Conformal Labeling’s selection result in regression task is highly sensitive to the choice of the tolerance parameter $\epsilon$. For instance, in sentiment analysis, setting $\epsilon=0.05$ yields an AI-labeled ratio of only $4.55\%$, whereas a slightly larger value $\epsilon=0.06$ increases the ratio to $47.84\%$. In practice, selecting an appropriate $\epsilon$ may require domain expertise or prior knowledge about the task.
Table 3: Performance of Conformal Labeling on regression tasks at $\alpha=0.1$. Results are shown for sentiment analysis (top) and protein folding (bottom) under different values of the tolerance parameter $\epsilon$. In all cases, Conformal Labeling controls the realized FDR below the target level.
Dataset Metric Method CL ($\epsilon=0.05$) CL ($\epsilon=0.06$) CL ($\epsilon=0.07$) Sentiment analysis FDR (%) 8.48% 8.60% 7.89% Power (%) 5.04% 53.14% 94.92% AI-labeled ratio 4.55% 47.84% 85.43% CL ($\epsilon=1$) CL ($\epsilon=4$) CL ($\epsilon=9$) Protein folding FDR (%) 9.67% 9.90% 9.04% Power (%) 27.24% 49.73% 97.90% AI-labeled ratio 10.06% 36.80% 88.00%

### v2

@@
-Selective labeling with false discovery rate control
+Model-agnostic Selective Labeling with Provable Statistical Guarantees
 Abstract
 Obtaining high-quality labels for large datasets is expensive, requiring massive annotations from human experts.
-While AI models offer a cost-effective alternative by predicting labels, their label quality is compromised by the unavoidable labeling errors.
-Existing methods mitigate this issue through selective labeling, where AI labels a subset and human labels the remainder.
-However, these methods lack theoretical guarantees on the quality of AI-assigned labels, often resulting in unacceptably high labeling error within the AI-labeled subset.
-To address this, we introduce Conformal Labeling, a novel method to identify instances where AI predictions can be provably trusted.
-This is achieved by controlling the false discovery rate (FDR), the proportion of incorrect labels within the selected subset.
-In particular, we construct a conformal $p$-value for each test instance by comparing AI models’ predicted confidence to those of calibration instances mislabeled by AI models.
-Then, we select test instances whose $p$-values are below a data-dependent threshold, certifying AI models’ predictions as trustworthy.
-We provide theoretical guarantees that Conformal Labeling controls the FDR below the nominal level, ensuring that a predefined fraction of AI-assigned labels is correct on average.
-Extensive experiments demonstrate that our method achieves tight FDR control with high power across various tasks, including image and text labeling, and LLM QA.
+While AI models offer a cost-effective alternative by predicting labels, their predictions inevitably contain errors.
+Existing methods mitigate this issue through selective labeling, where AI labels a subset and experts label the rest.
+However, these methods lack rigorous theoretical guarantees on the quality of AI-assigned labels, potentially leading to substantial labeling errors.
+To address this, we propose Conformal Labeling, a novel and model-agnostic method to identify a subset of unlabeled data with strict false discovery rate (FDR) control.
+In particular, we leverage a small labeled calibration set to construct conformal $p$-values for test data and estimate the accuracy of AI models.
+Then, we determine a data-dependent threshold for $p$-value selection, which automatically adapts to the estimated accuracy—higher accuracy permits a larger threshold.
+We provide theoretical guarantees that Conformal Labeling controls the FDR below the nominal level, ensuring that a predefined fraction of AI-assigned labels is correct in expectation.
+Extensive experiments across a wide range of models and datasets demonstrate that our method can achieve the target labeling quality with high power in classification and open-ended generation tasks.
+Selective Labeling, Uncertainty Quantification, Conformal Inference
 1 Introduction
 Large-scale, high-quality labeled data is crucial for the machine learning pipelines (Deng et al., 2009).
-While experts could provide high-quality labels for moderately sized datasets, the growing size of modern datasets has made this approach prohibitively expensive.
-AI models offer a cost-effective alternative by predicting labels, bypassing the need for human experts.
-However, AI models are prone to labeling error (Northcutt et al., 2021; Tan et al., 2024).
-For example, empirical evidence demonstrates that even state-of-the-art LLMs exhibit high labeling error when used for text annotation Baumann et al.
-(2025).
-The labeling error inherent to AI models significantly compromises their label quality, hindering the deployment of AI labeling for production.
-To balance the trade-off between labeling cost and error, selective labeling has been a promising solution (Geifman & El-Yaniv, 2017; Wang et al., 2023) by combining AI predictions with expert annotations.
-Prior work on selective labeling primarily designed heuristic methods (Wang et al., 2021; Bernhardt et al., 2022; Wang et al., 2024a) that rely on AI models to label high-confidence instances while deferring the rest to human experts.
-However, these methods do not provide any theoretical guarantees on label quality.
-To address this, a recent work (Candès et al., 2025) proposes probably approximately correct (PAC) labeling, which guarantees that the overall labeling error is controlled with high probability.
-They achieve the guarantee by counterbalancing the error of AI labels with the zero error of expert labels.
-While the guarantee is theoretically appealing, the labeling error of AI models can be unacceptably high, even reaching 100%, while the overall labeling error is controlled.
-These limitations motivate us to investigate how to provably guarantee the quality of AI-assigned labels.
-In this work, we propose Conformal Labeling, which formulates the labeling problem as multiple hypotheses testing.
-This method allows us to select a subset of the unlabeled dataset for AI labeling and ensure the quality of AI-assigned labels by providing a rigorous guarantee on the false discovery rate (FDR).
-In particular, given a labeled calibration dataset, we compute conformal $p$-values for test instances through a rank-based comparison of their predicted confidence against those of calibration instances where AI predictions are incorrect.
-This ensures that $p$-values for mislabeled instances stochastically dominate the uniform distribution on [0, 1], while $p$-values for correctly labeled instances are concentrated near zero, allowing selection procedures to distinguish between them easily.
-Then, we select test instances whose $p$-values are less than or equal to a data-driven threshold, which is delicately set with the calibration dataset to control FDR at the desired level.
-We further theoretically prove that Conformal Labeling achieves valid FDR control under mild assumptions, ensuring that the expected proportion of incorrect labels in the selected subset is below the desired level.
-While the label quality of existing methods depends heavily on model performance, Conformal Labeling guarantees label quality regardless of the underlying model’s performance by strictly controlling the FDR.
-To validate our method, we conduct extensive experiments on various labeling tasks, including image labeling (ImageNet (Deng et al., 2009), ImageNet-V2 (Recht et al., 2019)), text labeling (stance on global warming (Luo et al., 2021), and misinformation (Gabriel et al., 2022)), and LLM QA (MedMCQA (Pal et al., 2022), MMLU (Hendrycks et al., 2021), MMLU-Pro (Wang et al., 2024a)) tasks.
+While human experts can provide reliable annotations for moderately sized datasets, the rapid growth of modern datasets has rendered this approach prohibitively expensive.
+AI models offer a cost-effective alternative by predicting labels.
+However, AI models are prone to labeling error (Northcutt et al., 2021; Tan et al., 2024), with empirical evidence showing that state-of-the-art LLMs exhibit high error in text annotation tasks (Baumann et al., 2025).
+The prediction errors compromise label quality, hindering the deployment of AI labeling in real-world applications.
+To balance the trade-off between labeling cost and error, selective labeling has been a promising solution (Li et al., 2023; Wang et al., 2023) by combining AI predictions with expert annotations.
+Prior work on selective labeling primarily designed heuristic methods (Wang et al., 2021; Bernhardt et al., 2022; Wang et al., 2024a).
+These methods typically rely on model confidence, assigning high-confidence instances to AI models while deferring the remaining instances to human experts.
+However, these methods provide no theoretical guarantee on label quality.
+When the AI model is inaccurate or the confidence measure is unreliable, a substantial number of incorrect AI-assigned labels will be introduced, leading to a low-quality labeled dataset.
+Explicit accuracy guarantees directly address this issue by quantifying and controlling the quality of AI-assigned labels, ensuring that the resulting labeled dataset has a bounded error level pre-defined by users.
+This motivates us to design a model-agnostic method with provable accuracy guarantees on the AI-assigned labels.
+In this work, we propose Conformal Labeling, a novel method to identify a subset of unlabeled data with provable FDR control.
+In particular, we leverage a small labeled calibration set to construct conformal $p$-values for test data and estimate the accuracy of the AI model.
+Motivated by our analysis of the conservatism of the Benjamini–Hochberg (BH) procedure (Benjamini & Hochberg, 1995), we apply the BH procedure at an adjusted significance level derived from the estimated model accuracy, which enables effective FDR control while achieving high selection power.
+We provide theoretical guarantees that Conformal Labeling controls the FDR with distribution-free, finite-sample guarantees.
+While the error rate of AI labeling in current methods depends heavily on model performance, Conformal Labeling can control the desired labeling error in expectation regardless of the underlying model’s performance.
+We empirically validate our method through extensive experiments conducted on three labeling tasks, including image labeling (ImageNet (Deng et al., 2009), ImageNet-V2 (Recht et al., 2019)), LLM QA (MedMCQA (Pal et al., 2022), MMLU (Hendrycks et al., 2021), MMLU-Pro (Wang et al., 2024a)), and LLM open-ended generation (MMLU-Redux (Gema et al., 2025), MATH-500 (Lightman et al., 2023) MATH-L5 (Hendrycks et al., 2021), Zebra-Logic (Lin et al., 2025), and $\text{HumanEval}^{+}$ (Liu et al., 2023)) tasks.
 The results demonstrate that Conformal Labeling achieves high power with controlled FDR, indicating that AI models can label a large proportion of data with high quality.
-For example, Conformal Labeling can label 58.67% of the ImageNet dataset with ResNet-34 (He et al., 2016), while keeping the FDR below 10%.
-In comparison, a naive approach of using AI-assigned labels for the entire dataset results in a labeling error of over 25%.
-Moreover, through comprehensive ablation studies, we validate that Conformal Labeling is robust to the size of calibration datasets, and more powerful models enable better selection results.
+For example, Conformal Labeling can label 63.73% of Zebra-Logic with Qwen3-4B-Thinking-2507 (Yang et al., 2025), while controlling the FDR at 9.85%.
+In comparison, using AI-assigned labels for the entire dataset results in labeling errors of over 10%.
+Moreover, through comprehensive analyses, we validate that Conformal Labeling is robust to the size of calibration datasets, and the logits-based uncertainty score performs better than the verbalized uncertainty score.
 We summarize our contributions as follows:
-  • We propose Conformal Labeling, a novel method for identifying instances where AI predictions could be provably trusted.
+  • We propose Conformal Labeling, a novel method for identifying a subset where AI predictions could be provably trusted.
 Regardless of AI models’ performance, Conformal Labeling guarantees the quality of AI-assigned labels by strictly controlling the FDR.
-  • We theoretically prove that Conformal Labeling provides a strict quality guarantee for AI-assigned labels: it achieves valid FDR control, ensuring the expected proportion of incorrect labels is below a user-specified level.
-  • We empirically show that Conformal Labeling significantly reduces the labeling cost while tightly controlling the FDR, through extensive experiments conducted on image labeling, text labeling, and LLM QA tasks with various models.
+  • We theoretically prove that Conformal Labeling provides a strict quality guarantee for AI-assigned labels: it achieves FDR control, ensuring the expected portion of incorrect labels is below a user-specified level.
+  • We empirically show that Conformal Labeling can label a large portion of data with AI models while tightly controlling the FDR, through extensive experiments conducted on image labeling, LLM QA, and LLM open-ended generation tasks with various models.
 2 Preliminaries
 Problem setup.
-In this work, we study the problem of identifying instances where AI predictions can be provably trusted.
-Here, we give a formulation of multi-class classification as an example.
+We study the problem of identifying a subset where AI predictions can be provably trusted.
+Here, we give a formulation of multi-class classification as an example, while our framework also applies to open-ended generation.
 Let $\mathcal{X}$ denote the feature space and $\mathcal{Y}=\{1,\ldots,K\}$ denote the label space.
@@
 We denote the unseen ground-truth labels of instance $X_{j}$ as $Y_{j}$.
-Besides, we consider a pre-trained AI model $f:\mathcal{X}\to\mathbb{R}^{|\mathcal{Y}|}$ used to generate labels for the test dataset.
-For a given test instance $X$, the AI model predicts the label with the largest estimated probability $\hat{Y}=\arg\max_{y\in\mathcal{Y}}f_{y}(X)$, where $f_{y}(X)$ denotes the estimated class probability for class $y\in\mathcal{Y}$.
-Since AI models are typically prone to labeling errors, we aim to select a large subset from the test dataset $\mathcal{D}_{\mathrm{label}}$ to control the portion of incorrect labels.
-Formally, our goal is to identify the largest subset of indices $\mathcal{R}\subseteq\{1,\cdots,m\}$ that controls the false discovery rate (FDR), defined as below:
-$\text{FDR}=\mathbb{E}\left[\frac{|\mathcal{R}\cap\mathcal{H}_{0}|}{\max(|\mathcal{R}|,1)}\right],$
-where $H_{0}=\{j\in\{1,\cdots,m\}:Y_{j}\neq\hat{Y}_{j}\}$ is the set of indices with incorrect predictions.
+Besides, we consider a pre-trained AI model $f:\mathcal{X}\to\mathbb{R}^{|\mathcal{Y}|}$ used to predict labels for the test dataset.
+For a given test instance $X$, the AI model predicts the label with the largest estimated probability $\widehat{Y}=\arg\max_{y\in\mathcal{Y}}f_{y}(X)$, where $f_{y}(X)$ denotes the estimated class probability for class $y\in\mathcal{Y}$.
+Since AI models unavoidably make errors when used for labeling, we aim to select a large subset from the test dataset $\mathcal{D}_{\mathrm{label}}$ to control the portion of incorrect labels.
+Formally, our goal is to identify the largest subset of indices $\mathcal{R}\subseteq\{1,\cdots,m\}$ that controls the FDR, defined as below:
+$\mathrm{FDR}=\mathbb{E}\left[\frac{|\mathcal{R}\cap\mathcal{H}_{0}|}{\max(|\mathcal{R}|,1)}\right],$ (1)
+where $\mathcal{H}_{0}=\{j\in\{1,\cdots,m\}:Y_{j}\neq\widehat{Y}_{j}\}$ is the set of indices with incorrect predictions, and the expectation is taken over the randomness of the data.
 For notation shorthand, we denote $m]=\{1,\cdots,m\}$ in the following.
-The FDR metric measures the expected proportion of mislabeled samples within the AI-labeled subset, illustrating the quality of AI-assigned labels by explicitly bounding the fraction of incorrect labels.
-In addition to FDR control, we also expect AI models to label as many test instances as possible, which corresponds to maximizing power:
-$\text{Power}=\mathbb{E}\left[\frac{|\mathcal{R}\cap\mathcal{H}_{1}|}{\max(|\mathcal{H}_{1}|,1)}\right],$ (1)
+The FDR metric measures the expected proportion of mislabeled samples within the AI-labeled subset, illustrating the quality of AI-assigned labels by explicitly bounding the expected fraction of incorrect labels.
+[^1] The realized FDR might exceed the target level for a single run.
+In addition to FDR control, we also expect AI models to label as many test instances as possible correctly, which corresponds to maximizing power:
+$\mathrm{Power}=\mathbb{E}\left[\frac{|\mathcal{R}\cap\mathcal{H}_{1}|}{\max(|\mathcal{H}_{1}|,1)}\right],$ (2)
 […]
@@
 Besides, we assume that examples of the test and calibration datasets are both drawn i.i.d.
-from the joint distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$, a common setting in selective labeling (Jung et al., 2024; Candès et al., 2025).
+from the joint distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$.
+This assumption is realistic in our setting because we sample the calibration set from a large unlabeled dataset and leave the remaining data for testing.
+The process ensures that calibration data and test data are naturally i.i.d.
+In Appendix I.1, we test our proposed method against distribution shift to show that our proposed method is empirically robust to distribution shift.
 Selective labeling methods and their limitations.
@@
 Prior work on selective labeling primarily focused on heuristic methods.
-For example, some studies design collaborative annotation frameworks that combine expert labels with LLM labels to streamline the annotation process (Li et al., 2023; Kim et al., 2024).
-Others propose domain-specific methods, such as meta-learning strategies for medical image labeling (Vrabac et al., 2022), annotation frameworks for text data (Duan & Lalor, 2023), and human–AI collaborative systems for object detection (Zhang et al., 2025).
-Although these heuristic approaches effectively reduce annotation costs, they lack formal guarantees on the label quality, which can result in unreliable labels when AI models perform poorly.
-Probably approximately correct (PAC) labeling (Candès et al., 2025) addresses this limitation by providing a theoretical guarantee: the overall labeling error across the dataset is controlled with high probability.
-At its core, PAC labeling strategically collects zero-error expert labels for instances where the AI model exhibits the highest uncertainty, while relying on potentially imperfect AI predictions for more certain instances.
-This strategic allocation ensures that the dataset’s overall labeling error is small, as the zero-error expert annotations effectively counterbalance the error introduced by the AI-assigned labels.
-However, because the guarantee only applies at the aggregate dataset level, the subset labeled by AI may exhibit high labeling error.
-This limitation underscores a critical gap: existing selective labeling methods cannot ensure the quality of AI-assigned labels, hindering their reliable deployment in real-world applications.
-To address this, we explore methods to provably guarantee the quality of AI-assigned labels in selective labeling.
+For example, several works explore human-LLM collaboration for annotation, including uncertainty-guided work allocation (Li et al., 2023), verifier-based quality assessment that routes low-confidence samples to humans (Wang et al., 2024b), and unsupervised collaboration between LLMs and small models (Xiao et al., 2023).
+These methods balance labeling cost and quality by empirically selecting confidence thresholds to route samples between AI models and human annotators, yet provide no theoretical guarantees on the reliability of the AI-assigned labels.
+These methods implicitly assume well-calibrated confidence scores and sufficiently strong underlying model performance.
+However, commonly used confidence measures—including softmax probabilities from deep neural networks, logit-based confidence scores from LLMs, and LLM-generated verbalized confidence—are known to be poorly calibrated and overconfident (Guo et al., 2017; Chhikara, 2025), making them unreliable.
+Moreover, when the underlying model exhibits limited accuracy, even predictions assigned high confidence may be incorrect.
+Consequently, when the models are inaccurate or the confidence measure is unreliable, these methods will result in substantial labeling errors.
+These limitations motivate us to explore model-agnostic methods that can provably guarantee the quality of AI-assigned labels in selective labeling.
 3 Method
@@
 Our previous section shows that existing methods cannot guarantee the quality of AI-assigned labels.
-To address this limitation, we propose Conformal Labeling, which identifies instances where AI predictions can be provably trusted by controlling the FDR.
-Our approach is composed of three primary steps: quantifying uncertainty, constructing conformal $p$-values, and thresholding.
+To address this, we propose Conformal Labeling, a model-agnostic method to identify a subset where AI predictions can be provably trusted by controlling the FDR.
+Our method is composed of three primary steps: quantifying uncertainty, constructing conformal $p$-values, and thresholding.
 Uncertainty quantification.
-Our approach builds on a key insight: we should select instances where the model exhibits high confidence in its predictions.
-To quantify the model confidence, we define an uncertainty score $\mathcal{S}:\mathcal{X}\to\mathbb{R}$, where a higher value indicates greater model uncertainty.
-We note that in the conformal inference framework, this score function is also known as the non-conformity score function.
+Our method builds on a key insight: we should select instances where the model exhibits high confidence in its predictions.
+To quantify the model confidence, we define an uncertainty (non-conformity) score $\mathcal{S}:\mathcal{X}\to\mathbb{R}$, where a higher value indicates greater model uncertainty.
 For example, we employ $\mathcal{S}(X)=1-\max_{y\in\mathcal{Y}}f_{y}(X)$ as our uncertainty score function (Hendrycks & Gimpel, 2016).
@@
 To provide a statistical guarantee, we reformulate our problem as the following multiple hypothesis testing problem:
-$H_{j}^{0}:Y_{n+j}\neq\hat{Y}_{n+j}\quad\text{v.s.}\quad H_{j}^{1}:Y_{n+j}=\hat{Y}_{n+j},\quad\forall j=1,\cdots,m$
-Rejecting the null hypothesis $H_{j}^{0}$ indicates that $(X_{n+j},\hat{Y}_{n+j})$ should be included in the subset, as it is deemed to be classified correctly.
-To construct the selection subset, we employ conformal p-value, which builds upon conformal inference framework (Vovk et al., 1999; 2005).
-The underlying intuition is that: a test instance $X$ is likely a misclassification if its uncertainty score $\mathcal{S}(X)$ is generally larger than the scores of instances that are known to be misclassified.
+$H_{j}^{0}:\widehat{Y}_{n+j}\text{ is incorrect}\quad\text{v.s.}\quad H_{j}^{1}:\widehat{Y}_{n+j}\text{ is correct}.$
+The correctness of AI-assigned labels depends on the task definition.
+For image labeling tasks, correctness is defined as $\widehat{Y}_{n+j}=Y_{n+j}$.
+For open-ended verifiable tasks, correctness corresponds to the generated output being a correct solution.
+For alignment tasks, correctness means that the generated output aligns with human preferences.
+In the formulation, it is worth noting that rejecting the null hypothesis $H_{j}^{0}$ indicates that $(X_{n+j},\widehat{Y}_{n+j})$ should be included in the selection subset, as $\widehat{Y}_{n+j}$ is deemed to be correct.
+To construct the selection subset, we employ conformal p-value, which builds upon the conformal inference framework (Vovk et al., 1999, 2005).
+The underlying intuition is that: a test instance $X$ is likely to be misclassified if its uncertainty score $\mathcal{S}(X)$ is generally larger than the scores of instances that are known to be misclassified.
 Leveraging this idea, conformal $p$-value is computed through a rank-based comparison of $\mathcal{S}(X)$ against uncertainty scores of misclassified instances.
-Algorithm 1 Conformal Labeling
-1:Mislabeled Calibration set $\mathcal{D}^{0}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n_{0}}$, test instances $\{X_{n+j}\}_{j=1}^{m}$, pre-trained classifier $f$, calibration set size $n=|\mathcal{D}_{\mathrm{cal}}|$, FDR target $\alpha\in(0,1)$.
-2:# 1. Compute uncertainty scores
-3:for $i=1,\dots,n+m$ do
-4: Compute $\mathcal{S}_{i}:=1-\max_{y\in\mathcal{Y}}f_{y}(X_{i})$.
-5:end for
-6:# 2. Construct conformal p-values
-7:for $j=1,\dots,m$ do
-8: Construct $\hat{p}_{j}$ according to equation 2.
-9:end for
-10:# 3. Thresholding
-11:Compute $j^{*}=\max\left\{j:\hat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\}$, where $\hat{p}_{(j)}$ is the $j$-th smallest p-value.
-12:Output: $\mathcal{R}=\{j:\hat{p}_{j}\leq\hat{p}_{(j^{*})}\}$.
 Formally, for the calibration dataset $\mathcal{D}_{cal}$, we identify the subset $\mathcal{D}_{cal}^{0}\subseteq\mathcal{D}_{cal}$ where instances are misclassified by the AI model.
-For simplicity, we denote $\mathcal{D}_{cal}^{0}=\{(X_{i},Y_{i})\}_{i=1}^{n_{0}}$, and thus $Y_{i}\neq\hat{Y}_{i}$ for $i=1,\cdots,n_{0}$.
+For simplicity, we denote $\mathcal{D}_{cal}^{0}=\{(X_{i},Y_{i})\}_{i=1}^{n_{0}}$, and thus $\widehat{Y}_{i}$ is incorrect for $i=1,\cdots,n_{0}$.
 We compute the uncertainty scores for the entire dataset $\{(X_{i},Y_{i})\}_{i=1}^{n+m}$: $\mathcal{S}_{i}=1-\max_{y\in\mathcal{Y}}f_{y}(X_{i})$.
 Then, the conformal $p$-value for the instance $X_{n+j}$ is computed by
-$\hat{p}_{j}=\frac{\sum_{i=1}^{n_{0}}\mathbf{1}\{\mathcal{S}_{i}<\mathcal{S}_{n+j}\}+\left(1+\sum_{i=1}^{n_{0}}\mathbf{1}\{\mathcal{S}_{i}=\mathcal{S}_{n+j}\}\right)\cdot U_{j}}{n_{0}+1},$ (2)
+$\widehat{p}_{j}=\frac{\sum_{i=1}^{n_{0}}\mathbf{1}\{\mathcal{S}_{i}<\mathcal{S}_{n+j}\}+U_{j}\left(1+\sum_{i=1}^{n_{0}}\mathbf{1}\{\mathcal{S}_{i}=\mathcal{S}_{n+j}\}\right)}{n_{0}+1},$ (3)
 where $U_{j}\sim\mathrm{Uniform}[0,1]$ are i.i.d.
-uniform random variable to randomize over ties when $S_{n+j}$ equals some $S_{i}$, ensuring a continuous conformal $p$-value.
-Here, $\hat{p}_{j}$ quantifies how extreme the uncertainty of $X_{n+j}$ is compared to the scores of misclassified instances, with a small $\hat{p}_{j}$ providing strong statistical evidence for correct prediction.
-Standard results from conformal inference establish the validity of conformal $p$-value in practice: if the instance $X_{n+j}$ is misclassified, then the conformal $p$-value stochastically dominates the uniform distribution on $[0,1]$ (Bates et al., 2023; Jin & Candès, 2023): $\mathbb{P}\{\hat{p}_{j}\leq\alpha\mid H_{j}^{0}\text{ is true}\}\leq\alpha.$ This property indicates that the conformal $p$-values for misclassified instances are biased to be high, which allows us to set a threshold to flag potential correct instances, while controlling the FDR.
+uniform random variable to braek ties.
+Here, $\widehat{p}_{j}$ quantifies how extreme the uncertainty of $X_{n+j}$ is compared to the scores of mislabeled instances, with a small $\widehat{p}_{j}$ providing strong statistical evidence for correct prediction.
+The conformal $p$-value defined in (3) has a distribution-free finite-sample guarantee, which is formalized in the following proposition.
+The conservatism property indicates that the conformal $p$-values for misclassified instances are biased to be high as they stochastically dominate the uniform distribution on $[0,1]$.
+This allows us to set a threshold to flag potential correct instances, while controlling the FDR.
+Proposition 3.1.
+(from Vovk et al.
+(2005)).
+Given that the calibration sample $\{(X_{i},Y_{i})\}_{i=1}^{n}$ together with the $j$-th test sample $(X_{n+j},Y_{n+j})$ are i.i.d.
+for $j\in\{1,\dots,m\}$, conformal $p$-value $\widehat{p}_{j}$ defined in (3) is conservative:
+$\mathbb{P}\{\widehat{p}_{j}\leq\alpha\mid H_{j}^{0}\text{ is true}\}\leq\alpha,\quad\textnormal{for all }\alpha\in(0,1).$ (4)
 Thresholding.
-After obtaining conformal $p$-values for test instances, we apply a thresholding rule inspired by the Benjamini–Hochberg (BH) procedure (Benjamini & Hochberg, 1995) to select a maximal subset $\mathcal{R}$ for AI labeling while controlling the FDR at level $\alpha$.
-The key idea is that we gradually increase the acceptance threshold until including additional samples would risk exceeding the desired labeling error.
-In particular, let $p_{(1)}\leq\ldots\leq p_{(m)}$ denote the ordered statistics of the $p$-values; the rejection set of our selection procedure applied to the conformal $p$-values is $\mathcal{R}=\{j:\hat{p}_{j}\leq\hat{p}_{(j^{*})}\}$, where
-$j^{*}=\max\left\{j:\hat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\},$
+After obtaining conformal $p$-values for test instances, we can naively apply the Benjamini–Hochberg (BH) procedure (Benjamini & Hochberg, 1995) as in conformal selection (Jin & Candès, 2023) to select a subset $\mathcal{R}$ for AI labeling while controlling the FDR at level $\alpha$.
+Concretely, let $p_{(1)}\leq\ldots\leq p_{(m)}$ denote the ordered statistics of the $p$-values; the rejection set of the BH procedure applied to the conformal $p$-values is $\mathcal{R}=\{j:\widehat{p}_{j}\leq\widehat{p}_{(j^{*})}\}$, where
+$j^{*}=\max\left\{j:\widehat{p}_{(j)}\leq\frac{\alpha j}{m}\right\},$ (5)
 […]
-Theoretical results.
-We now provide a theoretical guarantee for Conformal Labeling.
-In the following theorem, we establish that Algorithm 1 controls the FDR below the desired level $\alpha$.
-This theorem ensures that AI predictions can be provably trusted in the selected subset, as the expected proportion of incorrect labels is controlled below $\alpha$.
-The proof is provided in Appendix B.2.
+To mitigate this conservatism issue, we estimate $\pi_{0}$ from the calibration dataset, building upon the intuition that when the calibration data and test data are i.i.d., the model should achieve comparable accuracy on both datasets.
+In particular, we estimate $\pi_{0}$ with $\frac{1+n_{0}}{1+n}$, and subsequently apply the BH procedure at level $\frac{1+n}{1+n_{0}}\alpha$.
+We summarize the complete procedure of Conformal Labeling in Algorithm 1, which combines all three steps described above.
+Algorithm 1 Conformal Labeling
+0: Calibration set $\mathcal{D}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, unlabeled test instances $\{X_{n+j}\}_{j=1}^{m}$, pre-trained model $f$, target FDR level $\alpha\in(0,1)$
+0: A selected subset $\mathcal{R}$ with FDR control
+1:// Identify mislabeled calibration data
+2: Identify $\mathcal{D}^{0}_{\mathrm{cal}}\subseteq\mathcal{D}_{\mathrm{cal}}$ with size $n_{0}$
+3:// Step 1: Compute uncertainty scores
+4:for $i=1$ to $n+m$ do
+5:  Compute uncertainty score $\mathcal{S}_{i}$
+6:end for
+7:// Step 2: Construct conformal $p$-values
+8:for $j=1$ to $m$ do
+9:  Construct conformal $p$-value $\widehat{p}_{j}$ according to Eq. (3)
+10:end for
+11:// Step 3: Thresholding
+12: Compute $j^{*}=\max\left\{j:\widehat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\}$, where $\widehat{p}_{(j)}$ is the $j$-th smallest $p$-value.
+13:Return $\mathcal{R}=\{j:\widehat{p}_{j}\leq\widehat{p}_{(j^{*})}\}$
+In Theorem 3.2, we establish that Conformal Labeling controls the FDR below the target level $\alpha$ even when it applies the BH procedure at level $\frac{1+n}{1+n_{0}}\alpha$.
+This result provides a formal guarantee that the AI-assigned labels within the selected subset are trustworthy, in the sense that the expected proportion of incorrect labels is bounded by $\alpha$.
+Because the BH procedure is applied at an elevated level, Conformal Labeling selects more samples than naively applying the BH procedure at level $\alpha$.
+Moreover, Conformal Labeling’s upper bound for FDR is tight: it converges to $\alpha$ as the calibration set size increases and equals $\alpha$ in the asymptotic limit.
+The proof is provided in Appendix A.2.
 Theorem 3.2.
-Suppose calibration samples $(X_{i},Y_{i})_{i=1}^{n}$ and test samples $(X_{n+j},Y_{n+j})_{j=1}^{m}$ are i.i.d.
+Suppose the calibration samples $(X_{i},Y_{i})_{i=1}^{n}$ and test samples $(X_{n+j},Y_{n+j})_{j=1}^{m}$ are i.i.d.
 Let $\alpha\in(0,1)$ be the target FDR level, and suppose the selection set $\mathcal{R}$ is determined by Algorithm 1 applied at the target FDR level $\alpha$.
 Define $p=\mathbb{E}[H_{j}^{0}]$, the probability that a test sample $(X_{n+j},Y_{n+j})$ is incorrectly predicted.
-Then, the FDR of the selection set $\mathcal{R}$ satisfies:
-$\text{FDR}\leq\left[1-(1-p)^{n+1}\right]\alpha\leq\alpha.$
-3.2 Choice of uncertainty score
-In the above analysis, we establish that Conformal Labeling controls the FDR below the desired level.
-However, this guarantee alone is insufficient: a trivial procedure that simply labels nothing would achieve a perfect FDR of 0, yet offer no practical value.
-This highlights the need to also evaluate the method’s statistical power, which measures the method’s ability to identify as many correctly labeled instances as possible (see Eq.
-(1)).
-As shown in prior work (Jin & Candès, 2023; Gui et al., 2024; Bai et al., 2025b), the statistical power of this method depends on the quality of the uncertainty score.
-In particular, a score that better separates correct from incorrect predictions directly increases statistical power (see Proposition 7 of Jin & Candès (2023)).
-To deliver practical recommendations, we empirically compare several uncertainty scores by visualizing their resulting $p$-value distributions and measuring the final statistical power employed with our method.
-We utilize a pre-trained ResNet-34 model on the ImageNet dataset, with three uncertainty scores: maximum softmax probability (MSP) (Hendrycks & Gimpel, 2016), energy score (Liu et al., 2020), and DOCTOR-$\alpha$ score ($D_{\alpha}$) (Granese et al., 2021).
-We give an overview of these score functions in Appendix C.
-The results in Figure 1 show that both MSP and $D_{\alpha}$ score provide a clear distinction between correct and incorrect predictions, thus achieving high statistical power.
-However, the energy score fails to provide this separation, leading to low power.
-Given the comparable power performance of MSP and $D_{\alpha}$ score, we will use the more computationally efficient MSP in our main experiments.
+Taking expectation over the randomness of the calibration and test data, the FDR of the selection set $\mathcal{R}$ satisfies:
+$\mathrm{FDR}\leq\left[1-(1-p)^{n+1}\right]\alpha\leq\alpha.$
 (a) MSP
@@
 However, the energy score fails to provide this separation, leading to low power.
+3.2 Choice of uncertainty score
+In the above analysis, we establish that Conformal Labeling controls the FDR below the desired level.
+However, this guarantee alone is insufficient: a trivial procedure that simply labels nothing would achieve a perfect FDR of 0, yet offer no practical value.
+This highlights the need to also evaluate the method’s statistical power, which measures the method’s ability to identify as many correctly labeled instances as possible (see Eq.
+(2)).
+As shown in prior work (Jin & Candès, 2023; Gui et al., 2024; Bai et al., 2025), the statistical power of this method depends on the quality of the uncertainty score.
+In particular, a score that better separates correct from incorrect predictions directly increases statistical power (see Proposition 7 of Jin & Candès (2023)).
+To deliver practical recommendations, we empirically compare several uncertainty scores by visualizing their resulting $p$-value distributions and measuring the final statistical power employed with our method in Figure 1.
+We utilize a pre-trained ResNet-34 model on the ImageNet dataset, with three uncertainty scores: maximum softmax probability (MSP) (Hendrycks & Gimpel, 2016), energy score (Liu et al., 2020), and DOCTOR-$\alpha$ score ($D_{\alpha}$) (Granese et al., 2021).
+We give an overview of these score functions in Appendix B.
+The results show that both MSP and $D_{\alpha}$ score provide a clear distinction between correct and incorrect predictions, thus achieving high statistical power.
+However, the energy score fails to provide this separation, leading to low power.
+Given the comparable power performance of MSP and $D_{\alpha}$ score, we will use the more computationally efficient MSP in our main experiments.
+In Appendix I.2, we provide a detailed comparison of the power and FDR achieved by these score functions.
 4 Experiments
-In this section, we evaluate the effectiveness of Conformal Labeling on image labeling, text labeling, and LLM QA tasks with various models.
-We find that it achieves tight FDR control and high power, indicating that AI models can label a large proportion of data with high quality.
-We also conduct comprehensive ablation studies to provide practical guidance for applying our method.
+In this section, we evaluate the effectiveness of Conformal Labeling on image labeling, LLM QA, and LLM open-ended generation tasks with various models.
+We find that it achieves tight FDR control and high power, indicating that AI models can label a large proportion of data with bounded error.
+In addition, we conduct additional analyses to provide practical guidance for applying our method.
 4.1 Experimental setup
-Tasks and datasets.
-We evaluate the effectiveness of Conformal Labeling across three labeling tasks, including image labeling, text labeling, and LLM question answering (QA).
-In appendix G, we also demonstrate how to extend Conformal Labeling to regression tasks.
-For the LLM QA task, the goal is to identify subsets of questions that LLMs can answer correctly.
-We employ common benchmark datasets for evaluations in each labeling task.
-For image classification, we use ImageNet (Deng et al., 2009) and its variant ImageNet-V2 datasets (Recht et al., 2019).
-For text labeling, we adopt two benchmark datasets.
-The first is Stance on Global Warming (Luo et al., 2021), which provides annotations ($Y_{i}\in\{\text{agree},\text{neutral},\text{disagree}\}$) to judge whether a headline agrees that global warming is a serious concern.
-The second is Misinformation (Gabriel et al., 2022), which contains binary annotations ($Y_{i}\in\{\text{misinfo},\text{real}\}$) for identifying whether a given text contains misinformation.
-For the LLM QA task, we evaluate our method on MedMCQA (Pal et al., 2022), MMLU (Hendrycks et al., 2021), and MMLU-Pro (Wang et al., 2024b) datasets.
-Table 1: Performance of Conformal Labeling on three labeling tasks.
-$\uparrow$ indicates that a larger value is better.
-We evaluate on Image Labeling (ImageNet, ImageNet-V2), text Labeling (Stance on global warming, Misinformation), and LLM QA (MedMCQA, MMLU) tasks.
-We report results for Conformal Labeling at $\alpha=0.1$ and compare Conformal Labeling against two baselines: (i) a naive approach of labeling instances whose uncertainty score $\mathcal{S}_{n+j}\leq 0.1$ with AI models, and (ii) labeling the entire dataset with AI models.
-The results show that Conformal Labeling consistently achieves tighter FDR control across all datasets and models compared with the baseline.
-Task Dataset Model Conformal Labeling ($\alpha$= 0.1) Naive ($\mathcal{S}\leq 0.1$) AI only
-FDR % Power %($\uparrow$) Ratio %($\uparrow$) FDR % Power %($\uparrow$) Ratio %($\uparrow$) Error %($\downarrow$)
-Image ImageNet ResNet-34 9.97 80.01 58.67 4.79 63.09 43.71 26.71
-DenseNet-161 9.99 85.03 65.56 5.58 72.08 52.98 22.89
-ResNeXt50 10.00 86.06 66.83 6.08 75.57 56.21 22.39
-CLIP-VIT-B/32 9.98 46.04 27.47 5.53 28.65 16.28 40.35
-ImageNet-V2 ResNet-34 10.00 61.99 37.87 7.25 56.03 33.15 39.03
-DenseNet-161 9.83 66.67 43.39 9.39 65.63 42.45 34.88
-ResNeXt50 9.95 67.86 44.75 10.17 68.71 45.38 34.08
-CLIP-VIT-B/32 9.96 34.56 18.10 7.86 25.83 13.18 47.78
-Text Stance Llama-3.1-8B-Instruct 9.77 19.02 9.42 24.43 54.94 31.39 52.04
-Qwen3-32B 9.56 10.70 7.25 14.42 16.64 11.11 36.52
-Qwen2.5-72B-Instruct 9.71 26.97 17.84 26.27 74.49 59.01 35.09
-Misinformation Llama-3.1-8B-Instruct 9.88 7.31 5.81 18.38 66.17 55.25 24.28
-Qwen3-32B 9.91 49.21 37.48 10.77 52.27 40.03 24.08
-Qwen2.5-72B-Instruct 9.58 44.36 34.81 17.81 87.13 74.72 21.69
-QA MedMCQA Llama-3.1-8B-Instruct 9.70 31.44 18.90 15.01 46.74 29.53 40.35
-Qwen3-32B 9.75 49.80 33.27 13.79 65.27 45.36 33.44
-Llama-3.1-70B-Instruct 9.95 69.67 49.59 4.52 48.92 32.79 28.90
-MMLU Llama-3.1-8B-Instruct 9.99 58.25 37.47 8.47 53.72 33.96 35.72
-Qwen3-32B 10.00 82.96 65.22 8.13 78.47 60.40 21.43
-Llama-3.1-70B-Instruct 9.96 88.20 72.10 4.18 67.94 52.17 18.24
+Datasets.
+We conduct experiments on image labeling, LLM QA, and LLM open-ended generation tasks.
+For image labeling, we use ImageNet (Deng et al., 2009) and ImageNet-V2 (Recht et al., 2019).
+For LLM QA task, we use MedMCQA (Pal et al., 2022), MMLU (Hendrycks et al., 2021), and MMLU-Pro (Wang et al., 2024c).
+For LLM open-ended generation task, we use MMLU-Redux (Gema et al., 2025), MATH-500 (Lightman et al., 2023), MATH-L5 (Hendrycks et al., 2021), Zebra-Logic (Lin et al., 2025), and $\text{HumanEval}^{+}$ (Liu et al., 2023) to evaluate Conformal Labeling’s performance in general, math reasoning, text reasoning and coding tasks respectively.
+We also show how to extend our method to regression tasks in Appendix H.
+Table 1: Annotation performance of Conformal Labeling on image Labeling and QA tasks.
+We compare Conformal Labeling against three baselines: (1) selection with guaranteed risk control (SGR), (2) FDR search heuristic method, and (3) labeling the entire dataset with AI models.
+We set $\alpha=0.1$ for Conformal Labeling and FDR search, and $\alpha=0.1$, $\epsilon=0.2$ for SGR.
+Bold numbers are superior results.
+Task Dataset Model Conformal Labeling SGR FDR search AI only
+FDR (%) Power (%) Ratio (%) FDR (%) Power (%) Ratio (%) FDR (%) Power (%) Ratio (%) Error (%)
+Image ImageNet ResNet-34 9.97 79.89 65.04 8.91 77.31 62.21 9.85 79.63 64.75 26.71
+DenseNet-161 9.92 84.89 72.68 8.92 82.68 70.00 9.83 84.71 72.45 22.89
+CLIP-VIT-B/32 9.84 45.63 30.21 8.17 39.66 25.79 9.77 45.41 30.04 40.35
+ImageNet-V2 ResNet-34 9.92 61.78 41.85 6.69 52.20 34.19 9.78 61.51 41.62 39.04
+DenseNet-161 9.98 67.07 48.57 6.82 54.18 37.98 9.84 66.67 48.21 34.88
+CLIP-VIT-B/32 9.99 35.52 20.73 8.17 6.59 3.67 9.51 33.61 19.55 47.78
+QA MedMCQA Llama-3.1-8B-Instruct 9.80 31.72 21.20 8.54 2.03 1.27 9.39 30.41 20.27 40.35
+Qwen3-32B 9.81 49.85 37.05 8.08 2.94 2.09 9.55 48.60 36.05 33.44
+Llama-3.1-70B-Instruct 9.86 69.37 54.83 6.43 44.34 33.67 10.58 71.15 56.67 28.90
+MMLU Llama-3.1-8B-Instruct 9.85 57.13 40.64 7.00 48.00 33.13 58.94 76.77 68.74 35.91
+Qwen3-32B 10.00 82.95 72.44 8.18 78.14 66.88 9.84 82.56 71.98 21.43
+Llama-3.1-70B-Instruct 9.97 88.25 80.17 8.21 84.00 74.85 9.90 88.10 79.98 18.24
+Table 2: Annotation performance of Conformal Labeling on open-ended generation tasks.
+We use the logits-based uncertainty score.
+We set $\alpha=0.05$ for Conformal Labeling and FDR search, and $\alpha=0.05$, $\epsilon=0.2$ for SGR.
+Bold numbers indicate superior results.
+Dataset Model Conformal Labeling SGR FDR search AI only
+FDR (%) Power (%) Ratio (%) FDR (%) Power (%) Ratio (%) FDR (%) Power (%) Ratio (%) Error (%)
+MMLU-Redux Qwen3-4B-Instruct-2507 4.84 33.62 28.46 43.17 9.06 7.71 4.22 29.87 25.11 20.17
+Qwen3-4B-Thinking-2507 4.95 24.64 21.46 42.56 4.08 3.61 5.85 23.77 20.62 20.79
+DeepSeek-R1-Distill-Qwen-32B 4.92 29.19 26.34 51.35 6.89 6.30 5.80 27.48 24.70 15.07
+MATH500 Qwen3-4B-Instruct-2507 4.93 86.73 83.88 3.18 52.83 50.72 4.88 84.94 82.10 8.20
+Qwen3-4B-Thinking-2507 4.07 99.80 99.78 2.50 57.25 57.11 4.01 99.69 99.67 4.00
+DeepSeek-R1-Distill-Qwen-32B 4.18 8.28 8.41 38.70 0.31 0.61 29.83 10.32 10.63 7.00
+MATH-L5 Qwen3-4B-Instruct-2507 4.84 69.54 66.90 0.55 2.90 2.85 5.14 68.89 66.25 8.88
+Qwen3-4B-Thinking-2507 4.64 48.33 48.45 0.35 1.96 1.95 4.72 57.85 57.94 5.27
+DeepSeek-R1-Distill-Qwen-32B 4.69 65.31 64.69 1.42 13.86 13.71 6.10 68.66 68.04 6.38
+Zebra-Logic Qwen3-4B-Instruct-2507 4.76 47.80 39.94 0.75 6.52 5.46 4.23 42.36 35.18 20.90
+Qwen3-4B-Thinking-2507 4.95 67.83 63.73 2.45 32.11 30.02 3.05 50.24 46.28 11.00
+DeepSeek-R1-Distill-Qwen-32B 4.78 0.01 0.02 69.33 0.03 0.03 67.70 0.05 0.32 31.40
+HumanEval+ Qwen3-4B-Instruct-2507 4.76 72.46 71.06 0.44 5.60 5.38 5.09 81.86 80.08 7.32
+Qwen3-4B-Thinking-2507 2.44 100.00 100.00 1.61 14.73 14.79 2.44 100.00 100.00 2.44
+DeepSeek-R1-Distill-Qwen-32B 0.61 100.00 100.00 0.64 39.48 39.49 0.62 100.00 100.00 0.61
 Models.
 We conduct extensive experiments on various open-sourced AI models.
-For image classification, we utilize three well-established deep image classifiers: ResNet-34 (He et al., 2016), DenseNet-161 (Huang et al., 2018), and ResNext50 (Xie et al., 2017).
-Additionally, we employ the Vision-Language Model CLIP (Radford et al., 2021), which is based on a Vision Transformer architecture (ViT-B/32) (Dosovitskiy et al., 2021).
-The above classifiers are provided by TorchVision (Paszke et al., 2019).
-For text labeling, we employ three LLMs: Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Qwen3-32B, Qwen2.5-72B-Instruct (Qwen et al., 2025).
-For LLM QA tasks, we employ five LLMs: Qwen3-8B (Yang et al., 2025), Qwen3-14B, Qwen3-32B, Llama-3.1-8B-Instruct, and Llama-3.1-70B-Instruct.
-The above LLMs are provided by Hugging Face.
+For image labeling, we use ResNet-34 (He et al., 2016), DenseNet-161 (Huang et al., 2018), and a Vision-Language Model CLIP (Radford et al., 2021), which is based on a pre-trained ViT-B/32 (Dosovitskiy et al., 2021).
+For LLM QA tasks, we employ Llama-3.1-8B-Instruct (Dubey et al., 2024), Qwen3-32B (Yang et al., 2025), and Llama-3.1-70B-Instruct.
+For LLM open-ended generation tasks, we use Qwen3-4B-Instruct-2507, Qwen3-4B-Thinking-2507, and DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025).
+The specific sampling temperature and other hyperparameters for LLMs are detailed in Appendix F.
+While the theoretical guarantee of Conformal Labeling is model-agnostic, we show in Appendix I.3 that higher AI model accuracy leads to a larger selection size.
+Uncertainty Scores.
+For the image labeling task, we use the MSP.
+For the LLM QA task, we also adopt MSP, where the estimated probability of each option is obtained by applying a softmax over the logits corresponding to the option tokens.
+For the LLM open-ended generation task, we adopt two complementary score functions that serve as analogues of MSP: a logits-based score for the white-box case and a verbalized score for the black-box case.
+We provide details for these score functions in Appendix E.
 Baselines and evaluation metrics.
-We evaluate Conformal Labeling with two baseline methods: using AI models to label test instances with uncertainty scores $\mathcal{S}_{n+j}\leq 0.1$, and applying AI predictions to the entire test dataset.
+We evaluate Conformal Labeling against four baseline methods: (1) an FDR-search heuristic baseline; (2) selection with guaranteed risk control (SGR) (Geifman & El-Yaniv, 2017); (3) labeling test instances with AI models when the uncertainty score satisfies $\mathcal{S}_{n+j}\leq\alpha$; and (4) applying AI predictions to the entire test dataset.
+We provide a detailed introduction to these baseline methods in Appendix G.
 We compare Conformal Labeling’s selection procedure against BH (Benjamini & Hochberg, 1995), Storey-BH (Storey, 2002), and Quantile-BH procedures (Benjamini et al., 2006).
-We evaluate the performance of our method and baseline using the following metrics: (1) FDR, the expected proportion of incorrect labels in the selected set $\mathcal{R}$; (2) Power, the proportion of correctly labeled instances selected; (3) AI-labeled ratio, the number of data labeled by the AI models divided by the combined size of the calibration and test datasets.
+We evaluate the performance of selective labeling using the following metrics: (1) FDR; (2) Power; (3) AI-labeled ratio, the proportion of data in the test dataset that is labeled by AI models.
 Implementation details.
-To ensure the reliability of our results, we repeat each experiment 1000 times and report the average result.
-We randomly select 10% of the data as the calibration dataset.
-For all the experiments, we use the maximum softmax probability (MSP) as the uncertainty score function.
-In the LLM QA task, we adopt the standard multiple-choice evaluation pipeline: given a question and candidate answers (e.g., A, B, C, or D), the model estimates the probability of each option by extracting the logits corresponding to the option tokens and applying a softmax transformation, with the predicted label taken as the option with the highest probability.
-For text labeling, we reformulate each sample into a multiple-choice format (e.g., “positive,” “negative,” or “neutral”), enabling the same probability-extraction procedure as in the LLM QA task.
+Each experiment is repeated 1000 times, and we report the mean FDR, power, and AI-labeled ratio.
 More details of implementation are provided in Appendix F.
 We provide the code for reproducing our main experiments in this anonymous repository.
-4.2 Experiments results
+4.2 Results
 Conformal Labeling achieves tight FDR control with high power.
-In table 1, we present the performance of Conformal Labeling against two baselines on three different labeling tasks: image labeling, text labeling, and LLM QA.
-A salient observation is that across all the labeling tasks and all the model architectures, Conformal Labeling successfully controls the FDR at or below the target FDR level.
-In comparison, both baseline methods lack FDR control, resulting in substantial labeling errors that compromise label quality when AI models are inaccurate.
-It is worth noting that the FDR is tightly controlled: for $\alpha=10\%$, most experiments yield FDRs below $9.9\%$, with the largest deviation at $9.56\%$.
-This tight FDR control directly leads to high selection power.
-For example, on MMLU with Qwen3-32B at $\alpha=10\%$, Conformal Labeling achieves a power of $82.96\%$, labeling $65.22\%$ of the dataset with the AI model.
-Overall, empirical results show that Conformal Labeling consistently achieves tight FDR control with high power across different datasets and models.
-Figure 2: Performance comparison of Conformal Labeling across models of varying accuracy.
-We employ Qwen3-8B, Qwen3-14B, and Qwen3-32B (model accuracy increases with parameter count) on MMLU and MMLU-Pro with $\alpha=0.1$.
-The results show that model with higher accuracy achieves greater power and AI-labeled ratio.
-Higher prediction accuracy enables better selection results.
-The performance of Conformal Labeling depends heavily on the underlying prediction accuracy, which is influenced by model capacity and dataset difficulty.
-We evaluate Conformal Labeling with Qwen3-8B, Qwen3-14B, and Qwen3-32B, whose increasing scales provide greater capacity.
-MMLU-Pro, more challenging than MMLU, results in lower prediction accuracy.
-Figure 2 presents the evaluation results.
-In all cases, the realized FDR remains below and close to $\alpha=0.1$.
-Higher accuracy—achieved with stronger models or easier datasets—boosts power and the AI-labeled ratio.
-Specifically, for any given model, performance is superior on MMLU compared to MMLU-Pro.
-Similarly, for each dataset, larger models yield greater power and AI-labeled ratios.
-Overall, higher prediction accuracy leads to better selection results by achieving higher power and AI-labeled ratio with controlled FDR.
+Table 1 reports the annotation performance of Conformal Labeling on image labeling and LLM QA tasks.
+Across all the benchmarks and model architectures, Conformal Labeling tightly controls the FDR while labeling a substantial proportion of data with AI models.
+For example, using Llama-3.1-70B-Instruct on MedMCQA, Conformal Labeling labels 54.83% of test instances with the AI model while controlling the FDR at 9.86%.
+In comparison, the FDR search baseline fails to control FDR, while SGR only labels 33.67% of test instances.
+Conformal Labeling improves the AI-labeled ratio by over 20% compared to SGR, while controlling the FDR.
+The tightest FDR control and highest power are also consistently observed in other datasets and models.
+Overall, Conformal Labeling achieves the best cost-accuracy tradeoff.
+Conformal Labeling outperforms baselines in the open-ended generation task.
+Table 2 reports the annotation performance of Conformal Labeling on open-ended generation task.
+The results show that Conformal Labeling reliably controls the FDR on open-ended generation tasks, whereas baselines may incur substantially violated FDR.
+For example, on MMLU-Redux with Qwen3-4B-Thinking-2507, Conformal Labeling controls the FDR at 4.95% while labeling 21.46% of instances, whereas SGR incurs an FDR of 42.56% and labels only 3.61% of data, and FDR search also exceeds the target FDR with an FDR of 5.85%.
+Besides, when baselines control the FDR below the target level, Conformal Labeling consistently achieves higher power than the baselines.
+Overall, Conformal Labeling achieves the best cost-accuracy tradeoff among all the compared methods.
+(a) MMLU-Redux
+(b) MATH-L5
+(c) Zebra-Logic
+Figure 2: Power of Conformal Labeling with differnet score functions.
+The experiments are conducted with Qwen3-4B-Instruct-2507 on three benchmarks.
+Shaded areas indicate standard deviations.
+We present the corresponding realized FDR in Figure 5.
+Logits-based uncertainty score outperforms verbalized uncertainty score in open-ended generation task.
+Figure 2 compares logits-based and verbalized score functions under varying target level $\alpha$.
+Under different score functions, Conformal Labeling consistently controls the FDR below the target level, while using logits-based score function results in higher power.
+For example, at $\alpha=0.07$, the logits-based score achieves over 50% power, whereas the verbalized score attains power of approximately 25%, corresponding to an absolute power improvement of over 25%.
+We also plot the reliability diagrams of these score functions in Figure 6, which shows that logits-based scores achieve lower expected calibration error (ECE) (Naeini et al., 2015) in most datasets, indicating superior reliability.
+We also plot the $p$-value distribution under these two score functions in Figure 7, which demonstrates that the logits-based uncertainty score provides a clearer separation between incorrect and correct labels.
+Together, these results demonstrate that the logits-based score provides more reliable uncertainty estimates than the verbalized score, leading to higher power.
+Conformal Labeling outperforms existing selection procedures.
+We conduct an ablation study on the selection procedure by fixing $p$-values and varying only the selection procedure.
+We compare Conformal Labeling’s selection procedure with three baselines: Storey-BH (Storey, 2002), Quantile-BH (Benjamini et al., 2006), and the standard BH procedure.
+Unlike Conformal Labeling that estimates $\pi_{0}$ from the calibration set, Storey-BH and Quantile-BH procedures estimate $\pi_{0}$ from the $p$-values.
+We provide a detailed introduction to these procedures in Appendix C.
+Besides, while Conformal Labeling’s selection procedure requires no hyperparameter tuning, Storey-BH and Quantile-BH involve hyperparameters that must be carefully specified.
+In this work, we set hyperparameters for these procedures using a bootstrap method (Storey, 2002) detailed in Appendix D.
+The comparison results in Table 3 show that our method can achieve much higher power than other selection procedures, with more precise FDR control.
+For instance, on MATH-L5 with Qwen3-4B-Instruct-2507, Conformal Labeling achieves a power of 94.25% at $\alpha=0.1$.
+In comparison, the BH, Storey-BH, and Quantile-BH procedures yield power of 0.09%, 61.42%, and 29.00%, respectively.
+Conformal Labeling’s selection procedure improves power by over 30% compared to these baselines.
+The highest power is also consistently observed in other datasets and models.
+In summary, Conformal Labeling’s selection procedure performs the best among all the compared selection procedures.
+Table 3: Comparison of different selection procedures on three labeling tasks.
+We compare Conformal Labeling’s selection procedure against three baselines: (1) BH procedure, (2) Storey-BH procedure, and (3) Quantile-BH procedure.
+We report results at $\alpha=0.1$ for all four selection procedures.
+Bold numbers are superior results.
+Task Dataset Model CL BH Storey-BH Quantile BH
+FDR % Power % FDR % Power (%) FDR (%) Power (%) FDR (%) Power (%)
+Image ImageNet ResNet-34 9.97 80.01 2.63 46.85 8.64 74.57 7.57 71.79
+DenseNet-161 9.99 85.03 2.27 44.55 8.34 81.17 7.33 78.40
+CLIP-VIT-B/32 9.98 46.04 4.08 21.70 8.52 41.02 8.05 39.23
+ImageNet-V2 ResNet-34 10.00 61.99 3.84 37.60 9.65 59.84 8.53 57.35
+DenseNet-161 9.83 66.67 3.39 28.39 9.62 65.45 8.44 61.66
+CLIP-VIT-B/32 9.96 34.56 4.85 14.43 9.20 32.25 8.35 28.52
+QA MedMCQA Llama-3.1-8B-Instruct 9.70 31.44 3.68 12.89 7.43 25.37 6.33 22.11
+Qwen3-32B 9.75 49.80 2.71 10.99 6.99 34.89 6.34 31.27
+Llama-3.1-70B-Instruct 9.95 69.67 2.79 29.67 8.48 64.09 6.74 58.06
+MMLU Llama-3.1-8B-Instruct 9.99 58.25 3.40 29.49 7.31 49.74 7.17 49.16
+Qwen3-32B 10.00 82.96 1.37 10.51 7.33 74.57 6.25 70.37
+Llama-3.1-70B-Instruct 9.96 88.20 1.61 18.74 6.95 80.08 5.96 76.77
+OE MMLU-Redux Qwen3-4B-Instruct-2507 9.96 67.62 1.85 5.52 7.50 50.52 5.15 36.72
+Qwen3-4B-Thinking-2507 9.88 66.37 1.59 1.71 4.93 26.86 3.77 15.58
+DeepSeek-R1-Distill-Qwen-32B 9.84 77.79 1.40 1.80 4.50 29.21 3.44 19.43
+MATH-L5 Qwen3-4B-Instruct-2507 7.97 94.25 0.99 0.09 5.38 61.42 3.13 29.00
+Qwen3-4B-Thinking-2507 5.16 95.69 0.68 0.02 1.69 5.66 0.91 1.61
+DeepSeek-R1-Distill-Qwen-32B 6.09 95.28 0.78 0.04 3.67 41.55 1.93 9.64
+Zebra-Logic Qwen3-4B-Instruct-2507 10.00 72.70 1.87 4.68 9.45 71.15 6.40 50.41
+Qwen3-4B-Thinking-2507 9.17 93.46 1.16 0.50 6.98 71.90 3.79 45.85
+DeepSeek-R1-Distill-Qwen-32B 8.65 0.14 3.69 0.01 2.80 0.11 3.02 0.01
 How many calibration samples are needed?
-The size of the calibration set plays a crucial role in constructing reliable conformal $p$-values.
-To study the effect of $|D_{\mathrm{cal}}|$ on FDR and power, in Figure 3, we label {5%, 10%, 20%} of the unlabeled dataset as the calibration dataset.
-Our results demonstrate that Conformal Labeling is robust to calibration set size: even with a 5% calibration ratio, the FDR remains controlled with low standard deviation.
+To study the effect of $|D_{\mathrm{cal}}|$ on FDR and power, in Figure 3, we label {5%, 10%, 20%} of the unlabeled dataset as the calibration set.
+Our results show that Conformal Labeling is robust to calibration set size: even with a 5% calibration ratio, the FDR remains controlled with low standard deviation.
 Increasing the calibration ratio reduces the variance of both FDR and power, although the improvement from 10% to 20% is negligible.
-Based on this trade-off between variance reduction and labeling cost, we use a 10% calibration ratio for all the experiments.
-In summary, while Conformal Labeling is robust to calibration set sizes, larger calibration sets reduce the variance of FDR and power, thereby enhancing selection stability.
-Figure 3: Performance of Conformal Labeling with varying calibration set sizes on the MedMCQA dataset.
-The top row corresponds to the results from Qwen3-8B and the bottom row to those from Qwen3-32B; each column corresponds to a value of calibration ratio.
-Shaded regions indicate one standard deviation around the average.
-The results show that a large calibration set slightly reduces the variance of FDR and power, leading to a more robust selection outcome.
-Overall, our method is robust to changes in the calibration set size.
-Figure 4: Albation study on the selection procedure.
-The ablation study is conducted on the MMLU dataset with Qwen3-32B.
-We substitute Conformal Labeling’s selection procedure with alternative selection procedures, including BH, Storey-BH, and Quantile-BH procedures.
-The figure shows the Conformal Labeling’s selection procedure consistently achieves the tightest FDR control and the highest power among all the selection procedures.
-Ablation study on the selection procedure.
-To validate the effectiveness of Conformal Labeling’s selection procedure, we conduct an ablation study substituting it with alternative selection procedures, including BH, Storey-BH (Storey, 2002), and Quantile-BH (Benjamini et al., 2006) procedures (see Appendix D for details).
-The hyperparameters for the Storey-BH and Quantile-BH procedures are chosen using a bootstrap method detailed in the Appendix E.
-An ideal selection procedure aims to maximize power while keeping the FDR below the desired level.
-The results of this ablation study are presented in Figure 4.
-While all the selection procedures successfully control the FDR below the desired level, Conformal Labeling’s selection procedure (red line) provides the tightest control, with realized FDR consistently closest to the desired level (dashed blue line).
-In comparison, alternative selection procedures are more conservative, especially at low target FDR levels, which reduces power.
-Overall, this ablation study highlights the superiority of Conformal Labeling’s selection procedure.
+Based on this trade-off between variance reduction and labeling cost, we use a 10% calibration ratio for image labeling and LLM QA tasks.
+Overall, while Conformal Labeling is robust to calibration set sizes, larger calibration sets reduce the variance of FDR and power.
+Figure 3: Performance of Conformal Labeling with varying calibration set sizes on MedMCQA with Qwen3-32B. The shaded areas represent standard deviations.
+Due to space constraints, we defer additional analyses to the appendix.
+In particular, Appendix I.5 compares our method against FDR search and selective prediction based on calibrated confidence baselines; Appendix I.6 examines whether the selected labels can be used to help train a model with improved accuracy on a synthetic dataset; and Appendix I.7 evaluates Conformal Labeling’s sensitivity to prompt designs in LLM QA tasks.
 5 Related work
-Selection with conformal $p$-values.
-Conformal $p$-values are widely studied for their distribution-free, model-agnostic properties in selection tasks.
-Existing methods fall into two main categories: conformal novelty detection (Bates et al., 2023; Bashari et al., 2023; Wu et al., 2025; Bashari et al., 2025; Lee et al., 2025; Huo et al., 2025) and conformal selection (Jin & Candès, 2023; Gui et al., 2024).
-Conformal novelty detection aims to identify out-of-distribution instances, while conformal selection aims to select data points that meet a specific quality criterion.
-For conformal novelty detection, recent advances primarily focus on improving the power of the selection procedure by using various forms of side information (Liang et al., 2024; Marandon et al., 2024; Zhao & Sun, 2025).
-Jin & Candès (2023) introduced the conformal selection framework for data selection in regression settings, which has since been extended to various settings: multivariate data selection (Bai et al., 2025b), online data selection (Xu & Ramdas, 2024; Liu et al., 2025), and human-in-the-loop adaptive data selection (Gui et al., 2025).
-Recent works have also applied CS to different tasks such as candidate screening (Lu et al., 2025), drug discovery (Jin & Candès, 2023; Bai et al., 2025a), and foundation model alignment (Gui et al., 2024).
-However, none of them has studied how to construct conformal $p$-values and design selection procedures in the selective labeling task.
-Selective labeling.
-Our method builds on the idea of selective labeling, which prioritizes the collection of expert labels for instances where the AI model exhibits uncertainty, and relies on the model’s prediction where it is confident.
-(Gu et al., 2012; Vrabac et al., 2022).
-Prior works on selective labeling have explored various heuristic methods, including collaborative annotation frameworks that integrate expert and LLM efforts (Li et al., 2023), and domain-specific applications in text, vision, and medicine (Kim et al., 2024; Vrabac et al., 2022; Duan & Lalor, 2023; Zhang et al., 2025).
-More recent research aims to provide theoretical foundations, with methods designed to control the overall labeling error (Candès et al., 2025) or to enable valid statistical inference (Zrnic & Candes, 2024; Gligorić et al., 2024).
-Our method is also closely related to selective prediction, where models are allowed to abstain from making a prediction when uncertain about the output (Geifman & El-Yaniv, 2017; Mozannar & Sontag, 2020; Yang et al., 2023).
-For LLMs, recent works have also studied how to teach the model not to predict when the model is uncertain (Kamath et al., 2020; Yoshikawa & Okazaki, 2023).
-These selective prediction methods also lack theoretical guarantees for the AI predictions.
-Distinct from prior methods, our work is the first to provide a guarantee for the quality of AI-assigned labels, thereby enabling the reliable deployment of selective labeling.
+Conformal Inference.
+Conformal inference (Vovk et al., 2005; Gazin et al., 2025; Humbert et al., 2025) is a statistical framework characterized by distribution-free and model-agnostic theoretical guarantees.
+In the literature, two particularly popular directions have emerged: conformal novelty detection (Bates et al., 2023; Bashari et al., 2023; Wu et al., 2025; Bashari et al., 2025; Lee et al., 2025) and conformal selection (Jin & Candès, 2023; Gui et al., 2024).
+The former focuses on identifying out-of-distribution instances, with recent advances focusing on enhancing selection power by incorporating various forms of side information (Liang et al., 2024; Marandon et al., 2024; Zhao & Sun, 2025).
+The latter aims to select candidates whose unobserved outcomes exceed user-specified values (Jin & Candès, 2023), which has been extended to multivariate data selection (Bai et al., 2025), online data selection (Xu & Ramdas, 2024; Liu et al., 2025), and human-in-the-loop adaptive data selection (Gui et al., 2025).
+Besides, some works have developed conformal inference techniques for classification tasks (Zhao & Su, 2023; Sun & Xia, 2025).
+For example, recent work (Zhao & Su, 2023) proposes controlling the overall error rate in binary and multi-class classification, while a concurrent work (Sun & Xia, 2025) extends this to control the general group-wise false discovery rate within a unified framework.
+In this work, we motivate and tailor the framework of conformal selection for selective labeling and support the application for open-ended generation tasks.
+Selective labeling and prediction.
+Selective labeling is an increasingly important topic for balancing the tradeoff between labeling cost and error (Gu et al., 2012; Vrabac et al., 2022).
+While prior work mainly explored heuristic approaches (Li et al., 2023; Duan & Lalor, 2023), a concurrent work (Candès et al., 2025) proposes approximately correct labeling.
+It guarantees that the overall labeling error is small with high probability by counterbalancing the error of AI labels with the zero error of expert labels.
+Our work considers a different objective: controlling the labeling error within the AI-labeled subset.
+Thus, these two methods address different error notions and are applicable in different settings.
+Notably, our method does not require expert annotations for samples on which the AI model abstains, whereas PAC labeling leverages such annotations to control the overall error.
+Some other work has also taken steps toward providing statistical guarantees for selective labeling (Zrnic & Candes, 2024; Gligorić et al., 2024), though these works focus on downstream inference rather than guaranteeing label quality.
+Our method is also related to selective prediction, where a model is allowed to abstain from making predictions when uncertain (El-Yaniv & Wiener, 2010; Mozannar & Sontag, 2020; Yang et al., 2023; Kamath et al., 2020; Yoshikawa & Okazaki, 2023).
+However, most of these methods cannot provide theoretical guarantees on prediction error.
+While works on selective prediction with risk control (El-Yaniv & Wiener, 2010; Geifman & El-Yaniv, 2017) can provide theoretical guarantees, they target a different error measure called marginal FDR, which is less preferred by prior researchers compared with FDR control.
+We provide a detailed discussion of the error measures in Appendix I.4.
 6 Conclusion
-In this work, we propose Conformal Labeling, a novel selective labeling method for identifying samples where AI predictions can be provably trusted.
-This is achieved by controlling the false discovery rate (FDR), which ensures that the expected fraction of incorrect labels in the selected subset is below a user-specified level.
-The key idea is to reformulate selective labeling as multiple hypothesis testing, which enables distinct theoretical guarantees and methodological advantages compared to prior approaches.
-In particular, we construct a conformal $p$-value for each test instance by comparing the AI model’s predicted confidence to those of mislabeled calibration instances.
-Then, we select all the test samples whose $p$-values are smaller than or equal to a data-dependent threshold.
-We theoretically prove that Conformal Labeling successfully controls the FDR under mild assumptions.
-Extensive experiments demonstrate that Conformal Labeling achieves tight FDR control and high power across various tasks, including image and text labeling, and LLM QA.
-We hope the insight from this work will inspire future research to explore reliable selective labeling.
-Appendix A Use of Large language model
-This paper uses large language models solely to polish specific sentences or paragraphs, without further use of LLMs for other purposes.
-Appendix B Technical proofs
-B.1 A lemma for proving Theorem 3.1
-Lemma B.1.
+In this paper, we introduce Conformal Labeling, a novel method for identifying a subset where AI predictions can be provably trusted.
+Specifically, we construct conformal $p$-values and estimate the accuracy of AI models using a small labeled calibration dataset.
+Then, we design a principled selection procedure that automatically adapts to the estimated accuracy, with higher accuracy permitting a larger threshold.
+We provide theoretical guarantees that Conformal Labeling controls the FDR below the nominal level under mild assumptions.
+Extensive experiments demonstrate that Conformal Labeling achieves tight FDR control and high power across various tasks, including image labeling, LLM QA, and LLM open-ended generation tasks.
+We hope that the insights from this work will inspire future research on rigorous methods for selective labeling.
+Appendix A Technical proofs
+A.1 Lemmas for proving Theorem 3.2
+Lemma A.1.
+(Benjamini & Yekutieli (2001), Theorem 1.2) If the joint distribution of the $p$-values is PRDS on the subset corresponding to true null hypotheses, applying the Benjamini–Hochberg procedure at level $\alpha$ guarantees
+$\text{FDR}\leq\frac{m_{0}}{m}\alpha,$
+where $m_{0}$ is the number of true null hypotheses.
+Lemma A.2.
 Under the assumptions of Theorem 3.2, where the calibration samples $(X_{i},Y_{i})_{i=1}^{n}$ and the test samples $(X_{n+j},Y_{n+j})_{j=1}^{m}$ are independently and identically distributed (i.i.d.), and conditional on $n_{0}$ (the number of true null hypotheses in the calibration set) and $m_{0}$ (the number of true null hypotheses in the test set), the expected false discovery proportion (FDP) satisfies
@@
 Proof.
-Suppose the first $n_{0}$ are true null samples in the calibration set and the first $m_{0}$ samples are true null samples in the test set.
-By the standard result in conformal inference (Vovk et al., 2005), we have
-$\mathbb{P}\{\hat{p}_{j}\leq\alpha\mid n_{0},m_{0}\}\leq\alpha,j=1,\ldots,m_{0}$
-Let $j^{*}$ denote the number of rejections from Algorithm 1, and recall the false discovery proportion (FDP) as
-$\text{FDP}=\frac{\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}}{\max(j^{*},1)}.$
-Since $\text{FDP}=0$ whenever $j^{*}=0$, we have
-$\mathbb{E}[\text{FDP}\mid n_{0},m_{0}]=$ $\mathbb{E}\!\left[\frac{1}{j^{*}}\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\middle|n_{0},m_{0},j^{*}>0\right]\mathbb{P}\left\{j^{*}>0\right\}$ (3)
-$\quad\quad+\mathbb{E}\!\left[\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\middle|n_{0},m_{0},j^{*}=0\right]\mathbb{P}\left\{j^{*}=0\right\}$
-$\leq$ $\mathbb{E}\!\left[\frac{1}{j^{*}}\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\middle|n_{0},m_{0},j^{*}>0\right]$
-$=$ $\mathbb{E}\!\left[\text{FDP}\;\middle|\;n_{0},m_{0},j^{*}>0\right].$
-By the super-uniform property of conformal $p$-values,
-$\mathbb{E}\!\left[\sum_{j=1}^{m_{0}}\mathbf{1}\!\left\{\hat{p}_{n+j}\leq\tfrac{1+n}{1+n_{0}}\cdot\tfrac{\alpha j^{*}}{m}\right\}\;\middle|\;n_{0},m_{0},j^{*}\right]\leq m_{0}\cdot\frac{1+n}{1+n_{0}}\cdot\frac{\alpha j^{*}}{m}.$
-Thus,
-$\mathbb{E}\!\left[\text{FDP}\;\middle|\;n_{0},m_{0},j^{*}>0\right]\leq\frac{1}{j^{*}}\cdot m_{0}\cdot\frac{1+n}{1+n_{0}}\cdot\frac{\alpha j^{*}}{m}=\frac{m_{0}}{m}\cdot\frac{1+n}{1+n_{0}}\alpha.$
-Combining with (3) yields
-$\mathbb{E}[\text{FDP}\mid n_{0},m_{0}]\leq\mathbb{E}\!\left[\text{FDP}\;\middle|\;n_{0},m_{0},j^{*}>0\right]\leq\frac{m_{0}}{m}\cdot\frac{1+n}{1+n_{0}}\alpha,$
-which completes the proof. ∎
-B.2 Proof of theorem 3.1
+Without loss of generality, assume that the first $n_{0}$ calibration samples and the first $m_{0}$ test samples correspond to true null hypotheses.
+By the standard result in conformal inference (Vovk et al., 2005) we have,
+[…]
+From results in (Bates et al., 2023), we know conformal $p$-values $(\widehat{p}_{1},\ldots,\widehat{p}_{m})$ are PRDS on the set of mislabeled data.
+Lemma A.1 hence guarantees that conditional on $n_{0}$ and $m_{0}$, applying Conformal Labeling at level $\alpha$ satisfies
+$\mathbb{E}[\text{FDP}\mid n_{0},m_{0}]\leq\frac{1+n}{1+n_{0}}\frac{m_{0}}{m}\alpha,$
+A.2 Proof of theorem 3.2
 Proof of theorem 3.2.
@@
 This completes the proof, establishing the desired bound on the FDR.
-Appendix C Overview of Different Uncertainty Score Functions
+Appendix B Overview of Different Uncertainty Score Functions
 In this section, we provide an overview of three representative uncertainty score functions that are widely used in misclassification detection: Maximum Softmax Probability (MSP) (Hendrycks & Gimpel, 2016), the energy-based score (Liu et al., 2020), and the DOCTOR-$\alpha$ score (Granese et al., 2021).
@@
 Compared to MSP, the DOCTOR-$\alpha$ score leverages information from the entire probability distribution rather than only the top prediction, making it a richer measure of uncertainty for misclassification detection.
-Appendix D BH Procedure and its Adaptive Variants
+Appendix C BH Procedure and its Adaptive Variants
 Consider testing $m$ null hypotheses $H_{0}^{1},\ldots,H_{0}^{m}$ based on their corresponding $p$-values $\{p_{1},p_{2},\ldots,p_{m}\}$.
@@
 If $\pi_{0}$ is small, the FDR control will be overly conservative.
+For example, if a classifier achieves a 80% accuracy in the unlabeled test dataset, then $\pi_{0}=0.2$—which leads to overly conservative FDR control.
 When $\pi_{0}$ is known, we can apply BH procedure at level $\frac{\alpha}{\pi_{0}}$ to close the gap.
@@
 Two most famous estimators are Storey-BH (Storey, 2002) and Quantile-BH (Benjamini et al., 2006):
-$\hat{\pi}_{0}^{Storey}(\lambda)=\frac{1+\sum_{i=1}^{m}\mathbf{1}\{p_{i}\geq\lambda\}}{m(1-\lambda)},\lambda\in(0,1)$
-$\hat{\pi}_{0}^{Quant}(k_{0})=\frac{m-k_{0}+1}{m(1-p_{(k_{0})})},k_{0}\in\{1,\dots,m\}$
+$\widehat{\pi}_{0}^{Storey}(\lambda)=\frac{1+\sum_{i=1}^{m}\mathbf{1}\{p_{i}\geq\lambda\}}{m(1-\lambda)},\lambda\in(0,1)$
+$\widehat{\pi}_{0}^{Quant}(k_{0})=\frac{m-k_{0}+1}{m(1-p_{(k_{0})})},k_{0}\in\{1,\dots,m\}$
 $\lambda$ and $k_{0}$ are hyperparameters determined by users.
-Appendix E Hyperparameter selection for BH adaptive variants
+Appendix D Hyperparameter selection for BH adaptive variants
 […]
@@
   5. 5.
-    Select $\hat{\lambda}=\operatorname*{arg\,min}_{\lambda\in R}\widehat{\mathrm{MSE}}(\lambda)$
+    Select $\widehat{\lambda}=\operatorname*{argmin}_{\lambda\in R}\widehat{\mathrm{MSE}}(\lambda)$
+Appendix E Uncertainty score functions for open-ended generation task
+Logits-based uncertainty score.
+In particular, let $y=(y_{1},y_{2},\ldots,y_{l})$ be a generated sequence of length $l$. We define the logits-based score as:
+$S_{\text{logits}}(x)=1-\frac{1}{l}\sum_{j=1}^{l}\mathbb{P}(y_{j}\mid y_{1},\ldots,y_{j-1},x)$
+where $\mathbb{P}(y_{j}\mid y_{1},\ldots,y_{j-1},x)$ is the conditional probability of token $y_{j}$.
+This score measures the average confidence of the model across all generated tokens.
+When the model is very confident about each token, the score is low.
+When the model is less confident, the score is high, indicating an example that may benefit from a more capable annotation source.
+Verbalized uncertainty score
+For the verbalized score, we first let the model explicitly state its self-reported confidence.
+The uncertainty score is then defined as one minus the self-reported confidence.
+In this work, we report the average confidence over 10 trials, and the prompts are listed in Table 4.
+Table 4: Prompt for the verbalized confidence scores.
+    System prompt: You are a reasoning assistant. For each question and proposed answer, you must estimate how likely the proposed answer is correct.
+    User prompt:
+    Question: {QUESTION}
+    Answer: {ANSWER}
+    Provide a probability (between 0.0 and 1.0) that your answer is correct. Only output the probability.
 Appendix F Implementation details
-Experiment details. We run our experiments on NVIDIA GeForce RTX 4090 and NVIDIA L40 GPU, and implement all methods by PyTorch and vLLM.
-Dataset details.
-For the LLM QA datasets (MedMCQA, MMLU, and MMLU-Pro), we adopt the same prompts as in Luo et al.
-(2025).
-For the other datasets, we design our own prompts following the style of LLM QA prompts, as summarized in Table 2.
-In text labeling and LLM QA tasks, we merge the calibration and test sets to form the unlabeled dataset, except for MedMCQA.
-For MedMCQA, because test labels are unavailable, we use the calibration set as the unlabeled dataset.
-For image labeling, we use the validation sets of ImageNet and ImageNet-V2 as the unlabeled datasets.
-Table 2: Prompts for different datasets
-   Dataset    Prompts    LLM QA    The following are multiple-choice questions.
-Give ONLY the correct option, no other words or explanation:    [Question] A: [Option 1] B: [Option 2] C: [Option 3] D: [Option 4] Answer: [Mask]        Stance on    global warming    You are given a statement about climate change.
-Determine the stance that a human would take towards this statement.
-Respond with ONLY the letter (A, B, or C) of the correct stance.
-Do not include any explanation.
-[Input Headline] A: agrees B: neutral C: disagrees Answer: [Mask]    Misinformation    You are a fact-checking assistant.
-Classify the following news headline as either real (A) or misinfo (B).
-Respond with ONLY the letter A or B.
-Do not include any explanation.
-Headline: [Input text] A: real B: misinfo Answer: [Mask]
-Appendix G Extension to Regression Tasks
+Hyperparameter settings of LLMs
+For the open-ended generation task, we configure the decoding parameters as follows.
+For Qwen/Qwen3-4B-Instruct-2507, we set Temperature=0.7, TopP=0.8, TopK=20, and MinP=0.
+For Qwen/Qwen3-4B-Thinking-2507, we set Temperature=0.6, TopP=0.95, TopK=20, and MinP=0.
+For DeepSeek/DeepSeek-R1-Distill-Qwen-32B, we set Temperature=0.6, TopP=0.95, and MinP=0.
+All experiments were run on NVIDIA GeForce RTX 4090, NVIDIA L40, and NVIDIA RTX Pro 6000.
+Dataset splitting strategy.
+For image labeling and LLM QA tasks, we randomly split 10\$ of the unlabeled dataset as the calibration dataset.
+For open-ended generation task, we use dataset-specific splitting strategy, which is detailed in Table LABEL:.
+Table 5: Details of dataset splitting settings for open-ended generation task.
+Dataset Dataset Type Dataset Size Split Setting Size
+MMLU-Redux General Task 3000 Calibration 500
+Test 2500
+MATH-500 Math Reasoning 500 Calibration 300
+Test 200
+MATH-L5 Math Reasoning 721 Calibration 300
+Test 421
+Zebra-Logic Text Reasoning 750 Calibration 300
+Test 450
+HumanEval+ Coding Task 164 Calibration 100
+Test 64
+Prompts for LLM QA task
+We adopt the same prompts as in Luo et al. (2025). The prompt is presented in Table 6.
+Table 6: Prompt for LLM QA task
+   Dataset    Prompt    LLM QA    The following are multiple-choice questions.
+Give ONLY the correct option, no other words or explanation:    [Question] A: [Option 1] B: [Option 2] C: [Option 3] D: [Option 4] Answer: [Mask]
+Appendix G Compared baselines
+We compare Conformal Labeling against four baseline methods: (1) an FDR-search heuristic baseline; (2) selection with guaranteed risk control (SGR) (Geifman & El-Yaniv, 2017); (3) labeling test instances with AI models when the uncertainty score satisfies $\mathcal{S}_{n+j}\leq\alpha$; and (4) applying AI predictions to the entire test dataset.
+FDR-search heuristic baseline selects the largest uncertainty threshold such that the empirical FDR of the calibration dataset does not surpass $\alpha$.
+SGR is a theoretically rigorous method that provides high-probability guarantees on marginal FDR control.
+It searches for the cutoff over a grid of confidence values, and then takes a union bound to achieve the overall error control.
+We provide a detailed introduction to the FDR-search heuristic baseline and SGR below.
+FDR-search heuristic baseline.
+Let $S(X)$ be a confidence function such as MSP. For a confidence threshold $t$, we define the empirical FDR on the calibration set as
+$\widehat{\mathrm{FDR}}(t)=\frac{\sum_{i=1}^{n}\mathbf{1}\{S(x_{i})\geq t,\;Y_{i}\neq\widehat{Y}_{i}\}}{\max\left\{\sum_{i=1}^{n}\mathbf{1}\{S(x_{i})\geq t\},\;1\right\}}$
+The threshold is selected by searching from the largest to the smallest confidence value.
+Specifically, we start from the smallest possible threshold and progressively increase $t$ until the empirical FDR on the calibration set falls below the target level $\alpha$:
+$t^{\ast}(\alpha)=\min\{\,t:\widehat{\mathrm{FDR}}_{\mathrm{cal}}(t)\leq\alpha\,\}.$
+Then a test instance is selected if its uncertainty score satisfies $\mathcal{S}(X_{i})\geq t^{\ast}(\alpha)$.
+Selection with guaranteed risk control (SGR).
+Let $f$ be a fixed classifier and let $c:\mathcal{X}\to\mathbb{R}$ be a confidence score, where larger values indicate more reliable predictions. For a threshold $\theta$, define the selection rule
+$g_{\theta}(x)=\mathbf{1}\{c(x)\geq\theta\},$
+so that predictions are made only on instances with confidence at least $\theta$. For a calibration set $\{(x_{i},y_{i})\}_{i=1}^{n}$, the empirical selective risk at threshold $\theta$ is
+$\widehat{R}(\theta)=\frac{\sum_{i=1}^{n}\mathbf{1}\{g_{\theta}(x_{i})=1,\;f(x_{i})\neq y_{i}\}}{\max}\left\{\sum_{i=1}^{n}\mathbf{1}\{g_{\theta}(x_{i})=1\},\,1\right\}.$
+The SGR algorithm searches over candidate thresholds (ordered from high to low confidence) and, for each threshold, computes a high-probability upper bound on the true selective risk based on the calibration data.
+The output threshold $\theta^{\ast}$ is the largest threshold whose risk bound does not exceed a target level $r^{\ast}$.
+At test time, an instance is selected if $c(X_{i})\geq\theta^{\ast}$.
+Theorem G.1 (Selective risk guarantee of SGR).
+Fix a target risk level $\alpha\in(0,1)$ and a confidence parameter $\delta\in(0,1)$.
+Let $\theta^{\ast}$ be the threshold returned by the SGR algorithm using a calibration set of size $n$.
+Then, with probability at least $1-\delta$ over the draw of the calibration set, the resulting selective classifier $(f,g_{\theta^{\ast}})$ satisfies
+$R(f,g_{\theta^{\ast}})\;\leq\;\alpha,$
+where $R(f,g_{\theta^{\ast}})$ denotes the true selective risk.
+In the labeling task, the risk function is defined as
+$R=\frac{\mathbb{E}\left[\sum_{j=1}^{m}\mathbf{1}\{Y_{n+j}\neq\widehat{Y}_{n+j}\text{ and }j\in\mathcal{R}\}\right]}{\mathbb{E}\left[\sum_{j=1}^{m}\mathbf{1}\{j\in\mathcal{R}\}\right]},$
+which is also known as the marginal false discovery rate (mFDR) in hypothesis testing.
+In the literature, FDR control is generally preferred over mFDR control.
+We provide a discussion of mFDR and FDR control in Appendix I.4.
+Appendix H Extension to Regression Tasks
 While our primary focus is on classification, Conformal Labeling can be naturally extended to regression settings.
-Consider a loss function L(Y,Y^)L($\hat{Y}$) that quantifies prediction error—for instance, the squared error $L(Y,\hat{Y})=(Y-\hat{Y})^{2}$—alongside a user-specified tolerance level $\epsilon$.
+Consider a loss function L(Y,Y^)L($\widehat{Y}$) that quantifies prediction error—for instance, the squared error $L(Y,\widehat{Y})=(Y-\widehat{Y})^{2}$—alongside a user-specified tolerance level $\epsilon$.
 For each test sample, we define the null and alternative hypotheses as:
-$H_{0}^{j}:L(Y_{n+j},\hat{Y}_{n+j})>\epsilon\quad\text{versus}\quad H_{1}^{j}:L(Y_{n+j},\hat{Y}_{n+j})\leq\epsilon.$
-This framework generalizes the classification setting, which corresponds to the special case where $L(Y,\hat{Y})=\mathbf{1}\{Y\neq\hat{Y}\}$ and $\epsilon=0$.
+$H_{0}^{j}:L(Y_{n+j},\widehat{Y}_{n+j})>\epsilon\quad\text{versus}\quad H_{1}^{j}:L(Y_{n+j},\widehat{Y}_{n+j})\leq\epsilon.$
+This framework generalizes the classification setting, which corresponds to the special case where $L(Y,\widehat{Y})=\mathbf{1}\{Y\neq\widehat{Y}\}$ and $\epsilon=0$.
 To apply Conformal Labeling in regression, we need an uncertainty score that reflects the model’s predictive uncertainty.
@@
 The complete procedure for Conformal Labeling in the regression task is outlined in Algorithm 2.
-Algorithm 2 Conformal Labeling Regression Tasks
-1:Calibration set $\mathcal{D}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$; test instances $\{X_{n+j}\}_{j=1}^{m}$; pre-trained predictor $f$; loss function $L$; loss threshold $\epsilon$; target FDR level $\alpha\in(0,1)$; nonconformity scores $\{S_{i}\}_{i=1}^{n+m}$
-2:Identify calibration samples that exceed the loss threshold: $\mathcal{D}_{\mathrm{cal}}^{0}=\{(X_{i},Y_{i}):L(Y_{i},\hat{Y}_{i})>\epsilon\}_{i=1}^{n}$, and set $n_{0}=|\mathcal{D}_{\mathrm{cal}}^{0}|$.
-3:for $j=1,\dots,m$ do
-4:  Compute the conformal $p$-value $\hat{p}_{j}$ according to equation 2.
-5:end for
-6:Apply the step-up procedure: $j^{*}=\max\left\{j:\hat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\}$, where $\hat{p}_{(j)}$ is the $j$-th smallest $p$-value.
-7:Output: The selected set $\mathcal{S}=\{j:\hat{p}_{j}\leq\hat{p}_{(j^{*})}\}$.
+Algorithm 2 Conformal Labeling for Regression Tasks
+0: Calibration set $\mathcal{D}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$; test instances $\{X_{n+j}\}_{j=1}^{m}$; pre-trained predictor $f$; loss function $L$; loss threshold $\epsilon$; target FDR level $\alpha\in(0,1)$; nonconformity scores $\{S_{i}\}_{i=1}^{n+m}$
+0: Selected set $\mathcal{S}$ with FDR control
+1:// Step 0: Identify calibration samples exceeding the loss threshold
+2: Set $\mathcal{D}_{\mathrm{cal}}^{0}=\{(X_{i},Y_{i}):L(Y_{i},\widehat{Y}_{i})>\epsilon\}_{i=1}^{n}$, and let $n_{0}=|\mathcal{D}_{\mathrm{cal}}^{0}|$.
+3:// Step 1: Construct conformal $p$-values
+4:for $j=1$ to $m$ do
+5:  Compute the conformal $p$-value $\widehat{p}_{j}$ according to Eq. (3).
+6:end for
+7:// Step 2: Step-up thresholding
+8: Compute $j^{*}=\max\left\{j:\widehat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\}$, where $\widehat{p}_{(j)}$ is the $j$-th smallest $p$-value.
+9:Return $\mathcal{S}=\{j:\widehat{p}_{j}\leq\widehat{p}_{(j^{*})}\}$.
 We evaluate Conformal Labeling on two regression problems: sentiment analysis with GPT-4o and protein structure prediction with AlphaFold.
-We use the data provided by Candès et al.
-(2025).
-For sentiment analysis, we prompt GPT-4o to output a prediction interval $a_{i},b_{i}]$ for each target $Y_{i}$, set the predicted value as $\hat{Y}_{i}=(a_{i}+b_{i})/2$, and use the interval length $U_{i}=b_{i}-a_{i}$ as the uncertainty score.
-For protein structure prediction, we take experimentally derived structures as ground truth and AlphaFold predictions as $\hat{Y}_{i}$.
+We use the data provided by (Candès et al., 2025).
+For sentiment analysis, we prompt GPT-4o to output a prediction interval $a_{i},b_{i}]$ for each target $Y_{i}$, set the predicted value as $\widehat{Y}_{i}=(a_{i}+b_{i})/2$, and use the interval length $U_{i}=b_{i}-a_{i}$ as the uncertainty score.
+For protein structure prediction, we take experimentally derived structures as ground truth and AlphaFold predictions as $\widehat{Y}_{i}$.
 Uncertainty scores are obtained from AlphaFold’s internal confidence measure, the average predicted local distance difference test (pLDDT).
@@
 In Table 7, we report the performance of Conformal Labeling on regression tasks with $\alpha=0.1$.
-In all cases, Conformal Labeling consistently controls the realized FDR below the target level.
-A key observation is that Conformal Labeling’s selection result in regression task is highly sensitive to the choice of the tolerance parameter $\epsilon$.
-For instance, in sentiment analysis, setting $\epsilon=0.05$ yields an AI-labeled ratio of only $4.55\%$, whereas a slightly larger value $\epsilon=0.06$ increases the ratio to $47.84\%$.
-In practice, selecting an appropriate $\epsilon$ may require domain expertise or prior knowledge about the task.
+In all cases, Conformal Labeling consistently controls the realized FDR below the target level, while a larger loss threshold $\epsilon$ leads to higher power.
 Table 7: Performance of Conformal Labeling on regression tasks at $\alpha=0.1$.
@@
 In all cases, Conformal Labeling controls the realized FDR below the target level.
-Dataset Metric Method CL ($\epsilon=0.05$) CL ($\epsilon=0.06$) CL ($\epsilon=0.07$) Sentiment analysis FDR (%) 8.48% 8.60% 7.89% Power (%) 5.04% 53.14% 94.92% AI-labeled ratio 4.55% 47.84% 85.43% CL ($\epsilon=1$) CL ($\epsilon=4$) CL ($\epsilon=9$) Protein folding FDR (%) 9.67% 9.90% 9.04% Power (%) 27.24% 49.73% 97.90% AI-labeled ratio 10.06% 36.80% 88.00%
+Dataset Metric Method CL ($\epsilon=0.05$) CL ($\epsilon=0.06$) CL ($\epsilon=0.07$) Sentiment analysis FDR (%) 8.48% 8.60% 7.89% Power (%) 5.04% 53.14% 94.92% CL ($\epsilon=1$) CL ($\epsilon=4$) CL ($\epsilon=9$) Protein folding FDR (%) 9.67% 9.90% 9.04% Power (%) 27.24% 49.73% 97.90%
+Appendix I Extensive Study
+I.1 Discussion of distribution shift
+We emphasize that the calibration set is sampled i.i.d.
+from a large unlabeled dataset (a transductive setting), leaving the remaining data for testing.
+Given a large unlabeled dataset, we first annotate a small subset as the calibration set, then apply conformal labeling to guarantee the quality of AI labeling.
+This process ensures that the calibration and test sets are naturally i.i.d., satisfying the standard assumptions of conformal labeling.
+In other words, Conformal Labeling generally does not encounter distribution shift in practice.
+While our method cannot provide theoretical guarantees under distribution shift, we add an experiment with ResNet34 on ImageNet and ImageNet-C Brightness to evaluate its empirical performance.
+In particular, we use ImageNet as the calibration set and ImageNet-C Brightness with varying severities as the testing set.
+We present the performance of our method across various testing sets in table 8.
+The results show that the realized power of our method is getting worse with a higher severity of distribution shift (may be due to the degraded accuracy), while the FDR is relatively insensitive.
+This demonstrates that our method is empirically robust to moderate distribution shift.
+Besides, we note that prior work (Jin & Candès, 2025) has investigated approaches for handling covariate shifts in conformal novelty detection and conformal selection.
+The same idea can be naturally incorporated into our Conformal Labeling procedure, enabling it to maintain false discovery rate control under covariate shift.
+We believe it can be an interesting direction for subsequent works if there are some realistic scenarios with distribution shift.
+Table 8: Performance of Conformal Labeling under varying severity levels of ImageNet-C Brightness corruption.
+”No shift” denotes that the test dataset also comes from ImageNet.
+Conformal Labeling maintains stable FDR control close to target levels while power decreases with increasing corruption severity.
+Test Dataset Accuracy (%) $\alpha=0.05$ $\alpha=0.1$
+FDR (%) Power (%) FDR (%) Power (%)
+No shift 73.29 5.00 72.27 9.85 79.62
+Severity 1 68.98 5.88 65.08 11.74 81.36
+Severity 2 67.17 5.87 63.55 11.77 80.12
+Severity 3 64.15 5.94 61.52 11.70 78.05
+Severity 4 59.30 5.62 57.31 11.67 75.14
+Severity 5 52.75 5.56 53.94 11.75 72.09
+I.2 Comparison of different score functions
+The choice of score function only affects the power but not the strict FDR control
+We present the results of three score functions on the ImageNet dataset in Table 9.
+While MSP, Energy, and $D_{\alpha}$ scores lead to vastly different power, all three successfully control the FDR below the target levels (5% and 10%).
+Specifically, MSP and $D_{\alpha}$ achieve high power (64% and 80% for $\alpha=5\%$ and $10\%$, respectively), whereas Energy yields significantly lower power (11.83% and 50.58%).
+This demonstrates that a poor score function can incur higher data collection costs by lowering power, but it does not compromise the statistical guarantee of FDR control.
+Table 9: Comparison of FDR control and power for MSP, Energy, and $D_{\alpha}$ score functions.
+The experiment is conducted with a ResNet-34 model on ImageNet.
+Bold indicates the highest power under the target FDR level.
+All methods maintain FDR close to target levels, while MSP and $D_{\alpha}$ achieve significantly higher power than the Energy score.
+Method Target FDR = 5% Target FDR = 10%
+FDR (%) Power (%) FDR (%) Power (%)
+MSP 4.97 63.87 9.97 80.01
+Energy 4.86 11.83 9.94 50.58
+$D_{\alpha}$ 4.95 63.80 9.92 79.74
+I.3 Impact of prediction accuracy
+Higher prediction accuracy enables better selection results.
+The performance of Conformal Labeling depends heavily on the underlying prediction accuracy, which is influenced by model capacity and dataset difficulty.
+We evaluate Conformal Labeling with Qwen3-8B, Qwen3-14B, and Qwen3-32B, whose increasing scales provide greater capacity.
+MMLU-Pro, more challenging than MMLU, results in lower prediction accuracy.
+Figure 4 presents the evaluation results.
+In all cases, the FDR remains below and close to $\alpha=0.1$.
+Higher accuracy—achieved with stronger models or easier datasets—boosts power and the AI-labeled ratio.
+Specifically, for any given model, performance is superior on MMLU compared to MMLU-Pro.
+Similarly, for each dataset, larger models yield greater power and AI-labeled ratios.
+Overall, higher prediction accuracy leads to better selection results by achieving higher power and AI-labeled ratio with controlled FDR.
+Figure 4: Performance comparison of Conformal Labeling across models of varying accuracy.
+We employ Qwen3-8B, Qwen3-14B, and Qwen3-32B (model accuracy increases with parameter count) on MMLU and MMLU-Pro with $\alpha=0.1$.
+The results show that model with higher accuracy achieves greater power and AI-labeled ratio.
+I.4 Discussion of mFDR control and FDR control.
+In the labeling task, the mFDR is defined as:
+$\mathrm{mFDR}=\frac{\mathbb{E}\left[\sum_{j=1}^{m}\mathbf{1}\{Y_{n+j}\neq\widehat{Y}_{n+j}\text{ and }j\in\mathcal{R}\}\right]}{\mathbb{E}\left[\sum_{j=1}^{m}\mathbf{1}\{j\in\mathcal{R}\}\right]},$
+while FDR is defined as
+$\mathrm{FDR}=\mathbb{E}\!\left[\frac{\sum_{j=1}^{m}\mathbf{1}\{Y_{n+j}\neq\widehat{Y}_{n+j}\text{ and }j\in\mathcal{R}\}}{\max\left\{\sum_{j=1}^{m}\mathbf{1}\{j\in\mathcal{R}\},\,1\right\}}\right].$
+There has been an exhaustive discussion of FDR and mFDR control.
+For example, Benjamini & Hochberg (1995) proposed the concept of FDR in favor of mFDR, because the latter is impossible to control in the global null case (Gui et al., 2024); Storey (2003) also pointed out that mFDR cannot be used for controlling the joint behavior of the numerator (number of false selections) and the denominator (number of selections).
+I.5 Comparison to selective prediction with calibrated confidences.
+Conformal Labeling achieves higher power than baselines under the same target FDR level.
+To compare our method with selective prediction with calibrated confidences, we conduct a new experiment on ImageNet using a calibrated ResNet-34 classifier.
+In particular, we use a holdout dataset to learn an optimal temperature parameter for temperature scaling, achieving an ECE of $2.22\%$.
+Given an error level $\alpha$, we compare Conformal Labeling with two heuristic baselines.
+(1) Confidence threshold.
+Given a target error level $\alpha$, an instance is selected if its calibrated maximum softmax probability (MSP) satisfies $\mathcal{S}(x)>1-\alpha$.
+(2) FDR search.
+We present the results of our method and two baselines in Table 10 below.
+The results show that Conformal Labeling consistently achieves the highest power and tightest FDR control among the three methods, showing that our method reliably increases the selected set size at the same target false discovery rate.
+Notably, we emphasize that the two heuristic baselines cannot provide any rigorous guarantee for the FDR control.
+We believe this demonstrates the benefits of our method compared to the baselines.
+Table 10: Comparison of different selection methods at target error levels $\alpha=0.05$ and $\alpha=0.10$.
+Conformal Labeling achieves the highest power and the tightest FDR control among the three methods.
+Method $\alpha=0.05$ $\alpha=0.10$
+FDR (%) Power (%) FDR (%) Power (%)
+Confidence threshold 2.80 57.23 3.94 66.60
+FDR search 4.51 63.11 9.65 75.90
+Conformal Labeling 5.00 72.27 9.97 79.91
+I.6 End-to-End Experiment with Conformal Labeling
+To assess whether Conformal Labeling can facilitate training a better downstream model, we conduct an end-to-end experiment that integrates label selection with classifier retraining.
+Specifically, we evaluate whether the samples selected by conformal labeling can effectively augment the training set and improve prediction performance.
+We adopt a synthetic dataset in this experiment to enable a controlled evaluation setting.
+In particular, synthetic data allows us to generate a sufficiently large number of i.i.d.
+samples from a fixed distribution, which allows us to better analyze the end-to-end behavior of conformal labeling.
+We generate a synthetic 10-class classification dataset with 27,000 samples using scikit-learn’s make_classification function. The dataset is randomly partitioned into three disjoint subsets:
+  • $D_{1}$: 2,000 samples used to train the initial labeling model;
+  • $D_{2}$: 20,000 samples used for AI labeling;
+  • $D_{3}$: 5,000 samples reserved for evaluation.
+We first train a base classifier on $D_{1}$.
+From $D_{2}$, we randomly select 2,000 labeled samples as calibration data and apply conformal labeling with a FDR of $\alpha=0.1$ to the remaining 18,000 unlabeled samples.
+This procedure yields an AI-labeled subset $D_{label}$.
+We repeat this process using three different base classifiers: KNN, MLP, and XGBoost.
+Table 11 reports the empirical false discovery proportion (FDP) and the number of selected samples for each classifier.
+In all cases, the observed FDP is well controlled below the target level, demonstrating the effectiveness of conformal labeling in selecting reliable pseudo-labeled data.
+Table 11: Selection results of conformal labeling with target FDR $\alpha=0.1$.
+Classifier FDP (%) Selection Size
+KNN ($k=5$) 9.72 6,142
+MLP (100, 25) 8.85 5,715
+XGBoost 9.56 4,006
+After obtaining $D_{\text{label}}$, we retrain each classifier using the combined dataset $D_{1}\cup D_{\text{label}}$.
+Table 12 compares the prediction accuracy on the held-out evaluation set $D_{3}$ between models trained only on $D_{1}$ and those trained on the augmented dataset.
+The results show that incorporating the selected samples consistently improves downstream performance.
+For example, when KNN is used as the base classifier, training on $D_{1}\cup D_{\text{label}}$ increases test accuracy from 64.92% to 69.60% (+4.68%).
+Similar gains are observed for MLP (+2.60%) and XGBoost (+0.26%).
+These findings indicate that conformal labeling enables effective utilization of unlabeled data to train more accurate models under some scenarios.
+Table 12: Prediction accuracy on $D_{3}$ with and without conformal-labeled samples.
+Classifier Training Set Accuracy (%) Improvement (%)
+KNN $D_{1}$ 64.92
+$D_{1}\cup D_{\text{label}}$ 69.60 +4.68
+MLP $D_{1}$ 65.68
+$D_{1}\cup D_{\text{label}}$ 68.28 +2.60
+XGBoost $D_{1}$ 59.80
+$D_{1}\cup D_{\text{label}}$ 60.06 +0.26
+I.7 Prompt Sensitivity Analysis
+To evaluate the sensitivity of our method to prompt design, we conduct experiments using Qwen3-32B on the MMLU dataset with three distinct prompts.
+In all experiments, the confidence measure is computed using a logits-based approach.
+The three prompts share identical question and answer content but differ in their phrasing and output format constraints.
+We describe each prompt in detail below.
+Prompt 1.
+The model is instructed to answer a multiple-choice question by outputting only the correct option without any additional text or explanation:
+The following are multi choice questions. Give ONLY the correct option, no other words or explanation:
+Question: {question}
+A: {choice1}
+B: {choice2}
+C: {choice3}
+D: {choice4}
+Answer:
+Prompt 2.
+The model is explicitly asked to respond with only the letter corresponding to the correct choice, again without providing explanations:
+You will be given multiple-choice questions. Respond with ONLY the letter of the correct choice. No explanations.
+Question: {question}
+A: {choice1}
+B: {choice2}
+C: {choice3}
+D: {choice4}
+Answer:
+Prompt 3.
+The model is prompted to output only the correct option, with a different structural layout that groups the answer choices under an “Options” header:
+Answer the following multiple-choice question. Output ONLY the correct option (A, B, C, etc.). No other text.
+Question: {question}
+Options:
+A: {choice1}
+B: {choice2}
+C: {choice3}
+D: {choice4}
+Correct option:
+Table 13 reports the performance of our method under the three prompt designs in terms of mean false discovery rate (FDR), mean power, and the testing error of the underlying language model.
+The results show that the mean FDR remains consistently close to the nominal level across all prompt designs, indicating that the FDR guarantee of our method is largely insensitive to prompt formulation.
+Besides, we notice that prompt 3 leads to lower power than the other two prompts.
+This is because the LLM achieves a higher testing error with this prompt.
+In summary, our method can be employed by LLM with different prompts.
+Table 13: Performance of our method under different prompt designs on the MMLU dataset.
+Prompt Mean FDR (%) Mean Power (%) Error (%)
+1 9.97 82.99 21.46
+2 9.94 82.58 21.48
+3 10.00 79.18 22.10
+Appendix J Additional experimental results
+In Figure 5, we present the realized FDR of Conformal Labeling under different target level on three benchmarks.
+The results show that ccross different target level, Conformal Labeling tightly controls the FDR in the open-ended generation tasks.
+(a) MMLU-Redux
+(b) MATH-L5
+(c) Zebra-Logic
+Figure 5: Empirical FDR of Conformal Labeling across target levels.
+Results are obtained using Qwen3-4B-Instruct-2507 on three benchmarks.
+The dashed line indicates the target FDR level $\alpha$.
+Shaded areas denote standard deviations.
+We present the reliability diagrams of Qwen3-4B-Instruct-2507 on MMLU-redux, MATH-L5, and Zebra-Logic in Figure 6.
+The results show that logits-based confidence achieves lower ECE than verbalized confidence in all datasets except for Zebra-Logic, indicating superior reliability of logits-based confidence.
+Figure 6: Reliability diagrams of Qwen3-4B-Instruct-2507 across three benchmark datasets (MMLU-Redux, MATH-L5, and Zebra-Logic).
+Each column corresponds to a dataset.
+The top row uses logits-based confidence, while the bottom row uses verbalized confidence.
+We present the $p$-value distribution of Qwen3-4B-Instruct-2507 using logits-based and verbalized confidence in Figure 7.
+The results show that the logits-based score provides a clearer separation between incorrect and correct predictions in most datasets.
+Figure 7: $p$-value distribution of Qwen3-4B-Instruct-2507 under different uncertainty score functions across three benchmark datasets.
+Each column corresponds to a dataset.
+The top row uses logits-based confidence, while the bottom row uses verbalized confidence.